A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.01

Title word cross-reference

1 [Woo43].

1939 [Ano40h]. 1940 [Ano41h]. 1941 [Ano42h].

75th [vO48].

Abundance [HD47]. Academic [Dea41]. Acclimation [Bre41]. Accompanying [Woo42]. Acid [PT47]. Acute [M’G41]. Administration [vOSW^{+}42]. Adults [Sla49]. advena [SS41a]. Advisory [Tri41]. Aerating [Bub40]. Aerial [HS48]. Africa [Du 48]. after [SM43]. Age [Car45a, DH47, HJ41, Job43, Job49, Sum48, VH49]. Agenda [Ano49a]. Ait. [SS41a]. Alabama [How42, SS40b]. Alaska [CGS49]. Albert [Raw45]. alcyon [SL49, SL49]. Alewife [RS45]. Algae [MHB48]. Algonquin [Ban41, FC48, Har41b]. Allotments [RS42]. alpenae [Job49]. alpinus [DM43]. Ambloplites [Bec41, Bec43, Hil42]. Ameiurus [RW40]. Amended [Ano40b, Ano41b, Ano42b, Ano43b, Ano47b]. American [Ano40h, Ano41h, Ano42h, Ano48a, Ano49i, Ano49j, vO47, AL45, Ano40b, Ano41b, Ano42b, Ano43b, Ano47b, vO48]. americana [Web43]. Ammunition [Deg45]. among [Dav47]. Anaesthesia [Fis43]. Analyses [Sho47b]. Analysis [LL49, Mot49, Vio45]. Andes [Tho40]. Anesthesia [GWS41]. Anglers’ [Cur40]. Angling [WL47]. Angora [NS42]. Animals [HWD48]. Anniversary [vO48]. Annual [Ano40h, Ano41h, Ano42h, Ano49a, Ano49v, vO47]. Apparatus [HWD48]. Appendages [Bel40b]. Application [Emb42, Fun49, HS48]. Appointment [Ano49b, Ano49c]. Appraisal [EWS^{+}42]. Aquarium [EH48]. Aquatic [Hog49a, SS42a]. Area [Els41, Lym43, LD45, SL43, Kin43]. Argentine [Tho40]. argus [Mar48]. Arkansas [Hog49a]. artedi [Car45b]. Arthur [Arm49]. Artificial [But41, Dav40, Rou47, Sur43b, Was48]. Associations [Rus41]. Atlantic [Bel40a, Bel40b, HR41, Rou47]. Attendance [Ano49o]. August [Ano42h, Ano42b, Ano43b, Ano47b]. Autumn [RL42]. Availability [Har41b]. Available [All42].

B [Phi49, Woo43]. Back [HS48, SF49]. Backwaters [Tar42]. Bantam [Web43]. Bar [Lym43]. Basin [Cot49]. Baskets [Car40]. Bass [Bec41, Bec43, Bro40, Du 48, Har41b, Hil42, Jon41, Raw45, SS43b, Sur40, Sur41, Sur43b, Sur45, Sur47, Sur49, Swi49, Vio43, Web48]. Bathymetric [Job43]. Bay [WFJ43, HJ41, Wal41]. Beaver [Spr41]. before [SM43]. Belted [SL49]. Better [Ano48b]. between [Bec48, MM45, Van42]. Biologists [Dea41]. Biology [NSH41]. Biotic [Hia47b]. Biotin [PB49]. Birch [LL49]. Black [Bro40, Har41b, Jon41, Kla40, Raw45, SS43b, Sur40, Sur41, Sur43b, Sur45, Sur47, Sur49, Swi49, Vio43]. Black-Head [Kla40]. Blackfly [CGS49]. Blackledge [TRW47]. Bloater [Sto47]. Bluegill [Kru49b, Leo40, SS40a]. Bluegills [MH45, SS41b, SS43b, Sur49, Swi49]. Blunt [Kla40]. Blunt-Nosed [Kla40]. Bottom [All42, How42, Lym43, LD45]. Bottom-Fauna [Lym43, LD45]. Bream [SS40a]. Brief [She49]. British [Mot40, Pri40]. Brook [BTH48, ESC^{+}41, Gut48, GS49a, GS49b, Kla41, Leo42, Phi47, PT47, Raw41, Smi41b, Smi47, TRW47]. Brown [EH48, GS49b, Hea43, Hey49, RL42, Sch45, SK48, TRW47]. Brunswick [Smi41a, Smi48]. Bullhead [RW40]. Bureau [Slo49]. Business [Ano49d, Ano49e]. By-Laws [Ano40b, Ano41b, Ano42b, Ano43b, Ano47b].

C [Jam43]. California [Ano40h, Bro40, CH40, Cro40, Cur40, NSH41, NS42, Sha40]. Can [Hig42]. Canada [Ano48b]. Cancellation [vO47]. Capacity [Pré41, SS40a]. Capture [Bre48]. Car [Mar48]. Carbon [Fis43, MM45]. Carbon-Dioxide [Fis43]. Carrying [SS40a]. Case [O’D45a]. Cass [LL49]. Catarrhal [M’G41]. Catch [Cur40, SM43]. Catches [Har41a, Mil45]. Catfish [BH48, Men45]. Caught [Sum48]. Cause [Dav49]. Cayuga [RW40]. Cedar [MH45]. Census [Mot49]. Central [Ric42, Nee49b, RL42]. cephalus [Hia47a]. Certain [Gut48, Tra41]. Certificate [Ano40a, Ano41a]. Certification [Ano42a, Ano43a, Ano47a]. Chains [Hia47a, Hia47b]. Changes [NS45, Tin42]. Channel [BH48]. Chanos [Hia47a]. Charrs [DM43]. Cherokee [LD45]. Chesapeake [Wal41]. Chinook [Cha40, PH45]. Chloride [Phi47]. Chub [Job43, Was48]. Cisco [Job49, Sto47]. Clair [Kru49a]. clarkii [Try43]. Classification [RS42]. Clear [Raw48]. Club [WL47]. clupeaformis [VH49]. Coarse [RG41]. Coast [Lin47, Smi40b]. Coefficients [Bec48]. Coho [Car40]. Cold [FR45]. Cold-Water [FR45]. Collecting [Bon40, Fun49, Has40, HZ41]. Columbia [Mot40, Pri40]. Column [O’D45a]. Columnaris [Dav49, FR45]. Combinations [Swi49]. Commercial [Bon41b, Car45a, Car45b, Car49, Lin47, Sum48, Tar45, Tri41]. Committee [Ano41f, Ano42f, Ano43g, Ano45e, Ano47g, Ano48e, Bag49, Dea49, Rei49, Tri41]. Committees [Ano40g, Ano40f, Ano41g, Ano42g, Ano43h, Ano45f, Ano47h, Ano48f, Ano49b, Ano49c, Ano49n, Ano49u, Ano49t, Ano49s, Ano49x, Ano49w, Bag49, Dea49, Rei49]. Common [Ano48b, RW40]. Comparison [All42, BTH48, Car40, Mil45, SM43]. Comparisons [Mot42]. Concentrations [Fis43]. Concluding [Ano49d, Ano49e]. Condition [Bec48, Kla41, Mot41, NS45]. Conditions [SK48]. Connecticut [TRW47, Web43]. Conservation [HS41, Luc41, Rus41, Whi49]. Considerations [Gri40]. Constitution [Ano40b, Ano41b, Ano42b, Ano43b, Ano47b]. Contained [RS42]. Content [Sur47]. Contents [LM41, MM45]. Control [Ano49f, CGS49, Eic49b, FB40, HT42, Hog49a, O’D45b, O’D47, Sha40, SS41a]. Controlled [Sch42]. Controlling [Fis47, SS42a]. Conversions [Bec48]. Cooperative [Rus41, Smi49]. Copper [Smi40a]. Coregonus [VH49]. Cornell [WL47]. Costia [Fis41]. Counts [BK42]. County [LL49, MH45, Sum48, WS43]. Covering [Try43]. Crecy [Smi48]. Creek [HS48, Nee40, SF49, Tre45, Was48, Sch45]. Creel [Mot49]. Creel-Census [Mot49]. Cresol [ESC^{+}41]. Criteria [Bel40b]. Critical [Vio45]. Croaker [Wal41]. Crustaceans [Lin47]. Crystal [Sch45]. Cultural [RS42]. Culture [Fis40, Rod47]. Current [Eic49a]. Cutthroat [Smi47, Try43]. Cycle [Hia47a, Hia47b]. Cytophaga [Dav49].

Dam [NSH41, Spr41]. Data [Mot49]. day [Mee43]. DDT [EH48, HS48, SF49]. Decline [Sum40, Van48]. Decomposition [MHB48]. Decreased [Cha40]. Deep [Ken41, Sto47, CA48]. Deep-Water [Sto47]. Deerskin [WS43]. Deficiency [Woo43]. Definition [Van49]. Deflectors [SCH49]. Density [Bec41, Bec43, Sch42, Sch45]. Depletion [TRW47, Van49]. Depth [Den48]. Derris [BB43a]. Design [Bre48]. Determined [Bel40a]. Determining [RS42]. Development [Cot49, Fel41, Nee49b, Pic49, Wal41]. Developments [Ano49f, HZ41]. Device [Bub40]. Diet [PH45]. Different [Kru49b, SK48]. Dioxide [Fis43]. Direct [Mot42]. Disease [DL41, FR45, Fis47, O’D47, Wol41]. Diseases [Dav47]. Disinfection [O’D47]. Dissolved [CW49]. Distribution [Den48, Gri40, HD47, Job43, Job49, Wal41]. dolomieu [Har41b, Sur41, Sur43b, Sur43a, Web48]. Dosage [GS49a]. Dosages [GS49b]. Drag [CH40]. Drag-net [CH40]. Drainage [DM43]. Drugs [Gut48]. during [Hig42, Van47].

Early [Ken41, Wal41]. Eastern [SL49, Kla41, Raw41, Smi47, Sur41]. Eating [Sim40]. Ectoparasites [Emb42]. Education [Whi49]. Effect [DF41, ESC^{+}41, Esc43, Law41, Mot41, Phi47, Rou47, Sch42, SS40a, Spr41, Try43, Wri45]. Effects [Cha40, CGS49, Cot49, Hau49, HS48, Hog41, SCH49, Sla49, Sur45]. Efficient [Rou42]. Effort [Sch42]. Egg [BK42, Sim40]. Egg-Eating [Sim40]. Eggs [Bel40b, Car40, Hey49, SB40, Tin42, Web48]. election [Ano40c, Ano41c, Ano42c, Ano43c, Ano47d]. Electric [Hau49]. Electrical [Fun49, Has40, HZ41]. Elimination [Ano49f, Bal48b]. Elops [Hia47a]. Emergency [Hig42]. England [HR41, RS45]. Enteritis [M’G41]. Epidemics [Dav49]. Erie [Van42, Van48, VH49]. Esox [Car42, Kru49a]. Especial [AL45]. Especially [Gut48]. Estimating [Emb40, Raw49]. Ether [GWS41]. Europe [Nee49b]. Evaluating [HWD48]. Evaluation [Fis40, PH45]. Examples [CS45]. Excised [Sla49]. Exotic [Gri40]. Experiment [Bal47, BB43a, RG41, SB40]. Experimental [Bjo40, DF41, Raw45, Sur49]. Experiments [Bal48b, Bel40a, FB40, Rou42, SV40, Swi49, WL47]. Exploited [Ric49]. Expressing [Hil48]. External [Sur43a].

Facilities [Dea41]. Factor [Van48]. Factors [Bec48, SM45]. Failure [Raw48]. Fall [She47, Van47]. Family [AL45]. fario [SK48]. Farm [Sur49]. Fast [SK48]. Fasting [Law41]. Fauna [Lym43, Spr41, LD45]. Faunas [All42]. Features [Pic49]. Fecundity [Kat48]. Feeding [But41, Hia47a, Leo40, Leo42, LL49, MH45]. Fertility [Hey49]. Fertilization [Nee49b, Smi48]. Fertilized [How42, MM45, SS41b]. Fertilizer [SS42a]. Fertilizers [Hog49a, SS40a, SS43a, Sur45]. Field [Rou42]. Fin [Arm49, Hea43]. Fingerling [DF41, SK48]. Fingerlings [Arm49, Law41, Leo42, M’G41, PH45, Sla49]. Fins [Sla49]. First [Ano42h]. Fish [All42, Ano49f, Arm49, Bal47, Bal48a, Bal48b, Ban41, BB43a, BB43b, Bub40, CS45, CA48, CW49, Cot49, Dav49, Den48, Du 48, Eic49b, Emb40, Emb42, Fun49, Gre41, GJSL41, Har41a, HWD48, Has40, HZ41, Hey49, Hia47a, Hia47b, Hil48, Hof41, HS48, Hog41, How42, HSR^{+}42, Hun48, Ken41, Lan42, Mee42, Mee43, MM45, Mil45, Nee40, NS42, O’D43, Raw49, RG41, RS42, Rod47, RK45, SL49, Sha47, Sho47a, Smi40a, SV40, SS43a, SJH49, Sur47, SS40b, SS42b, Tar42, Tar45, Van49, Vio45, Wri45]. Fish-Cultural [RS42]. Fish-Food [BB43a, HS48]. Fish-Stocking [Lan42]. Fisheries [Ano40b, Ano40h, Ano41b, Ano41h, Ano42b, Ano42h, Ano43b, Ano47b, Ano48a, Ano49i, Ano49j, Bon41b, Car45a, Car45b, Car49, Cro40, Hig42, RS45, Tri41, vO47, vO48]. Fishermen [Sum48]. Fishery [CH40, Dea41, Eic49a, Fel41, FC48, HS41, Hil49, Luc41, Men45, Rus41, Smi49, Sto47, Sum40, Tar45, Whi49, vOSW^{+}42, Tri41]. Fishes [Ano48b, Bec48, Bec49, Dav47, Deg45, FB40, FR45, Gag43, Gri40, Lin47, MHB48, MA45, O’D45a, Ric42, Ric49, SF49, Tho40, Tra41, Van48, Woo42]. Fishing [Cro40, Esc43, HT42, Sch42, TM43, Wri45]. Five [Hil42, Mee42, Tra41]. Five-Year [Tra41]. flavescens [HJ41, MH45]. Flesh [PH45]. Flood [Ano49f]. Flood-Control [Ano49f]. Florida [Mee42]. Fluctuations [Car45a, Car45b, Tra41]. Fly [WL47]. following [Bal48a, Bec41, Bec43, NS45]. fontinalis [ESC^{+}41, Gut48, GS49a]. Food [All42, BH48, Bon41a, BB43a, DF41, Fel41, Hia47a, Hia47b, Hig42, HS48, Leo42, RW40, RL42, Sur41, Vio45, Web43]. Food-Chains [Hia47a, Hia47b]. Foods [Nee40]. Forage [Kla40]. Ford [Bal48a, Leo40]. Forest [Cha43]. Formalin [FB40]. formation [Sla49]. Forms [AL45]. Formulations [SF49]. Found [Sur43a]. Four [Kla41]. Fourth [vO47]. Francisco [Ano40h]. Fresh [Ell42, Ric49]. Fresh-Water [Ric49, Ell42]. Freshets [Hun48]. Fry [But41, Car40, Kru49b, Van42]. Fungus [DL41]. Further [Bec43, HZ41, Leo40, She47, Wol41]. Furunculosis [Gut48, GS49a, O’D47].

Gain [Mee43]. gairdnerii [Mot42, PM41, Sum48]. Game [Eic49b]. Gars [Bon41a]. Gear [CH40]. Geochemical [Sho47b]. Gill [Sto47]. Gills [Sur43a]. Girard [Ran42]. given [Emb40]. Glen [Jam43]. Gmelin [Web43]. Gonad [BK42]. Gorbuscha [Pri40]. Gordon [Ano49o]. Granules [Gag43]. Gravel [Car40, SB40]. Grayling [Leo40]. Great [Raw49, Van48, WFJ43]. Greater [Vio45]. Grounds [HD47]. Growing [Jon41]. Growth [Arm49, Bec41, Bec43, Bec49, Car45a, Car45b, DH47, DF41, HJ41, Hil42, Job43, Job49, Kru49b, LM41, Mar48, Mar41, Mot42, NS45, RW40, Sch45, SK48, SS41b, Sum48, Try43, VH49, Vio43]. Growths [Sur45]. Gulf [Bel40a]. Gyrodactyloid [Du 48]. Gyrodactylus [Emb42].

Habitats [Mil45]. Habits [BH48, DM43, Hia47a, Leo40, Leo42, LL49, MH45, Sim40]. Haddock [Rou42]. Hampshire [WFJ43]. Hard [Sur45]. Hard-Water [Sur45]. Harmful [Hau49]. Hatcheries [O’D47, RML45]. Hatchery [Arm49, BTH48, Car40, Du 48, Emb42, Fis47, Har41a, Hey49, Hog49a, Kru49b, Pré41, SK48, She47, Try43, Woo43]. Hatchery-Raised [Har41a]. Hatchery-Reared [BTH48, She47]. Hatching [SB40]. Hawaiian [Hia47a, Hia47b]. Head [Kla40]. Heavy [MHB48]. Herring [Kat48]. High [Fis43, Hog41, NS42]. History [Car42, MA45, PM41, She49]. Hoop [Mil45, SM43]. Hoop-Net [Mil45, SM43]. Houghton [Car42]. hoyi [Sto47]. Huro [SS43b]. Huron [Van47, Hil49]. Hybognathus [Ran42]. Hybrid [Ric48]. Hydrodictyon [O’D45b].

Ictalurus [BH48]. II [Hia47b]. Impoundment [Esc43, Lym43, SM43, LD45]. Impoundments [Ell42, Tar45]. Improvements [Cla48]. inch [SS45]. Incorporation [Ano40a, Ano41a, Ano42a, Ano43a, Ano47a]. Increased [Bec41, Bec43]. Increasing [Mot41, Pré41]. Incubated [Car40]. Incubation [Web48]. Indiana [RML45]. Indices [Moy49]. Individual [RS42]. Infectious [Dav47, Fis47]. Infestation [Kla40]. Infesting [Emb42]. Influence [Sur45]. Influencing [SM45]. Information [Luc41]. Initial [Raw48]. Inland [Du 48, Hey49]. Inorganic [SS40a]. Insect [Spr41]. Insecticides [CGS49]. Inspection [Hof41]. Institutional [Dea41]. Instructions [Ano48a, Ano49i, Ano49j]. Intensities [Kru49b]. Intensity [Esc43, HT42, TM43]. Interaction [Hia47b]. International [Dea49]. Interpretation [Sho47b]. Introduced [MA45]. Introduction [Gri40, MA45, Raw45, Raw48]. Investigation [Has41]. Investigations [MM49]. Iowa [BH48, Ros49].

Jasper [Raw41]. Jonkershoek [Hey49, Du 48]. June [Ano40h, Ano40b].

Karluk [DM43]. Kept [Mar48]. Key [AL45]. King [Sla49]. Kingfisher [SL49]. kisutch [SB40]. Kiyi [DH47, HD47, DH47, HD47]. Known [Ano48b]. Koelz [DH47, HD47, Job43, Job49].

L. [Car42]. Laboratory [Sho47a]. Lacépède [Har41b]. lacustris [BH48]. Lagoon [Nee40]. Lake [Arm49, Bal47, Bal48a, Car42, CA48, FC48, Hil49, Hol49, Ken41, Kin43, Leo40, LL49, MH45, Mot40, Mot41, Moy49, NS42, RW40, Raw48, Ray42, RG41, Ros49, SV40, Smi41a, Smi48, Tra41, VH49, Web43, Wri45, BB43a, BB43b, Car45a, Car45b, Car49, DH47, HD47, Hil49, Job43, Job49, Kru49a, Raw49, Ric42, SV40, Sto47, Sum40, Van42, Van48, VH49, Woo42]. Lake-Management [Hol49]. Lakes [Bal48b, Ban41, CW49, EWS^{+}42, Har41b, Hil42, Mee42, NS42, O’D43, Raw45, SH41, She47, Van47, Van48]. Lamprey [MM49, She49]. Large [Hau49]. Largemouth [Du 48, Jon41, SS43b, Sur40, Sur49, Swi49, Vio43]. Lawrence [Bel40a]. Laws [Ano40b, Ano41b, Ano42b, Ano43b, Ano47b]. Leach [Jam43]. Legal [Sch42, SH41, She47, Smi41b, WS43]. Legal-Sized [Sch42, She47, WS43]. Length [Bec48, Car45b, Jon41, Kru49a, Roy42, Smi41b]. Lengths [Bec48, Hil48, Tra41]. Lepomis [Leo40, MH45, SS43b]. LeSueur [Bjo40, Mar41, RW40]. Leucichthys [Car45b, DH47, HD47, Job43, Job49, Sto47]. Liberation [GJSL41]. Life [Car42, PM41]. Limited [SL43]. Ling [Bjo40, Mar41]. Linnaeus [SL49]. List [Ano40c, Ano41c, Ano42c, Ano43c, Ano47d, Ano48b, Ano49k, Ano49l]. Little [Ric49]. Little-Exploited [Ric49]. Live [Mar48]. Lobsters [Mar48]. Localized [Eic49b]. Lockers [HS41]. Longjaw [Job49]. Lonoke [Hog49a]. lota [Mar41, Bjo40, Mar41]. Louis [Ano42h]. Louisiana [Vio43]. Lower [TM43]. lucius [Car42]. m [Kru49a]. machnata [Hia47a]. macrochirus [Leo40, MH45, SS43b]. Maculosa [Mar41, Bjo40]. Madison [BK42]. Maine [RS45]. Maintaining [Rou47]. Maintenance [O’D47]. make [Mot42]. Maligne [Raw41]. malma [DM43]. Management [Cha43, Dav40, Eic49b, EWS^{+}42, Hol49, Kin43, NS42, RS45, SL49, SS42b, Tar42, Vio45]. Manitoba [Raw48]. Manuscripts [Ano48a, Ano49i, Ano49j]. Marine [Cro40, Lin47]. marinus [MM49]. Mark [RK45]. Marked [Arm49, Bal48a, NS42, SH41, Sla49]. Market [Luc41]. Marketing [Rus41]. Maryland [HT42]. Maskinonge [Els41]. Masquinongy [Kru49a]. Materials [SS43a]. Maturity [Car45a]. Measure [Mee43]. Measurement [GJSL41, Rod47, TM43, Tin42]. Measurements [BK42, Sha47]. Measuring [Emb42]. Meeting [Ano40e, Ano40h, Ano41e, Ano41h, Ano42e, Ano42h, Ano43e, Ano45c, Ano47f, Ano48d, Ano49o, Ano49p, Ano49q, vO47, Ano49a, Ano49v]. Megaceryle [SL49]. Members [Ano40c, Ano41c, Ano42c, Ano43c, Ano47d, Ano49k, Ano49l, Ano49o]. membership [Ano40c, Ano41c, Ano42c, Ano43c, Ano47d]. Memoriam [Ano45a, Ano47c, Ano49g, Ano49h]. Method [Emb40, Emb42, Fun49, Has40, HZ41, Pré41]. Methods [Bon40, Hil48, RS42, Rod47]. Mexico [Lin47]. Michigan [Bal47, Bal48a, Bal48b, Bec48, Bec49, Car42, CS45, CA48, Cla48, CW49, DH47, HD47, Job43, Job49, Leo40, LL49, MH45, SH41, SL43, She47, She49, SCH49, SV40, Smi41b, Tra41, Van47]. Micropogon [Wal41]. micropteri [Sur40]. Micropterus [Har41b, Sur41, Sur43b, Sur43a, Web48]. Migrating [Bre48]. Migration [Bel40a, Ken41, Ray42]. Milkfish [Hia47a]. Mining [Smi40b]. Minnesota [Car45a, Car45b, Car49, Hof41, SM45, SS45, SJH49]. Minnow [Ran42]. Minnows [Kla40, Swi49]. Mississippi [Smi49]. Missouri [Ano42h, Cot49, Pic49]. Mitchill [Car45a, HJ41, MH45, Van47, VH49]. Modified [Tes41]. Moines [BH48]. Molluscs [Lin47]. Molting [Mar48]. Montana [BK42, Leo40]. montanus [Leo40]. Moonlight [Wri45]. mordax [Van47, WFJ43]. Morone [Web43]. Mortality [Arm49, CW49, DF41, MHB48, NS45, Ric49, Van47, Woo42]. Most [Rou42]. Mountain [Raw48]. Movement [Sch45]. Mr. [Ano49o]. Mugil [Hia47a]. Mullet [Hia47a]. Multiple [HSR^{+}42]. Muskellunge [Kru49a].

Names [Ano48b]. National [Cha43, Hig42, Raw41, Raw45]. Native [SF49]. Natural [Leo42, Pri40, Rou47, Smi47, Sur43b]. Navigation [Ano49f]. Neascus [Kla40]. nebulosus [RW40, RW40]. necatrix [Fis41]. Net [Mil45, SM43, CH40]. Nevada [MA45]. News [Luc41]. Niacin [PB49]. Ninth [Ano40h]. Nitrogen [MM45, Sur47]. None [BB43a]. Norris [Jon41]. North [Kat48]. Northeastern [Hil42]. Northern [Car42, HR40, Ric42, Smi41b, O’D43]. Nosed [Kla40]. Notes [Fis41, NSH41]. nuchalis [Ran42]. Number [Bel40b, Emb40, Emb42]. Numbers [RS42, Tra41]. Nuphar [SS41a].

Observations [Bjo40, Car42, Leo40, Leo42, Nee40, Sim40, Smi48, Sur43b, Wol41]. Officers [Ano40d, Ano41d, Ano42d, Ano43d, Ano43f, Ano45b, Ano45d, Ano47e, Ano48c, Ano49m, Ano49n, Ano49r, Hun49, vO47]. Oil [SF49]. Oncorhynchus [Pri40, SB40]. Ontario [Ano41h, Arm49, Har41a, Har41b, MM49, Sto47]. Opening [Hog49b, vO49]. Operation [Bre48]. Orderly [Rus41]. Oregon [Gri40, Hol49, Sum48]. Organic [SS40a, SS43a]. Organisms [BB43a, HS48, How42, Leo42]. Osmerus [Van47, WFJ43]. Ossification [O’D45a]. Other [Kla40, SV40]. Our [Fel41]. Oxygen [Cha40, CW49, Has41, Woo42]. Oxygenating [HWD48]. Oyster [Bon40].

Pacific [Kat48, Smi40b]. Palatability [BTH48]. Pantothenic [PT47]. Panulirus [Mar48]. Parasite [Du 48, Sur40]. Parasites [Ban41, HR40, Ric42]. Park [Har41b, Raw45, Raw48, Ban41, FC48, Raw41]. Part [Hia47a, Hia47b]. Particular [Emb42, Sto47]. Parts [Kat48]. Paul [Mot40]. Penaeidae [AL45]. Pennsylvania [Tre45]. Penobscot [Rou47]. Perca [HJ41, MH45]. Percentages [SS43b]. Perch [Car45a, HJ41, LM41, MH45, Web43]. Period [Tra41]. Peritrichous [Sur43a]. Petromyzon [MM49]. Phenomenal [Vio43]. Photograph [Ano49o]. Physiological [Phi47]. Pickerel [HR40]. Pike [Car42, Car45a, HR40]. Pike-Perch [Car45a]. Pikeperch [Ros49, SM45]. Pink [Pri40]. Pisgah [Cha43]. Place [HSR^{+}42, Bag49]. Placer [Smi40b]. Places [Ano40e, Ano41e, Ano42e, Ano43e, Ano45c, Ano47f, Ano48d, Ano49p, Ano49q]. Plankton [SS40a]. Plant [Deg45, Sur45]. Planted [NS42, RL42, SK48, SS43b, SS45]. Planting [Bre41, NS45, Smi41b, WS43]. Plantings [Har41a, SH41, She47, Tre45, Van42]. Plants [Hog49a, SS42a]. Poisoning [Bal48a, Gre41]. Poisons [Smi40a]. Policies [EWS^{+}42, Lan42, Slo49]. Pollution [Ano49f, Rei49, Sha40]. Pond [Nee49b]. Ponds [EWS^{+}42, Hia47a, Hia47b, Hog49a, How42, Kru49b, Mee43, MM45, O’D45b, Ran42, Ric48, SS40a, SS41a, SS41b, SS42a, SS43a, SS43b, SM45, Sur45, Sur47, Sur49, SS40b, SS42b, Swi49, Was48]. Population [Bal47, Bal48a, Bec41, Bec43, BB43b, CA48, Car45a, Car45b, LM41, O’D43, Ros49, Sch42, Sch45, SL43, TRW47]. Populations [Bal48b, Mee42, Ric49, Sho47a, SJH49, SS42b, Tar42]. Port [Arm49]. Possibilities [Bon41b, Tar45]. Potter [Smi41a]. Pounder [Hia47a]. Practices [Vio45]. Pre [Lym43, LD45]. Pre-Impoundment [Lym43, LD45]. Predicting [Den48]. Preliminary [Bjo40, Sim40, Smi48]. Preparing [Ano48a, Ano49i, Ano49j]. Present [Leo42]. Presentation [Ano49o]. President [Hog49b, vO49]. Presidents’ [Ano40e, Ano41e, Ano42e, Ano43e, Ano45c, Ano47f, Ano48d, Ano49p, Ano49q]. Prince [Raw45]. Private [Har41a]. Probable [Sur45]. Problem [She49]. Problems [NSH41, NS42, Tar45]. Production [HJ41, HSR^{+}42, Mee43, MM45, Mot40, Raw49, RML45, SS40a, SM45, Sur45, Sur47, SS40b, Van42, Vio45]. Productivity [Moy49]. Products [HS41, RS42]. Program [HSR^{+}42, Slo49, Pic49]. Programs [Eic49a, Lan42, vOSW^{+}42]. Progression [Web43]. Projector [Tes41]. Prolonged [FB40]. Promotion [Bon41b]. Propagation [Bro40, Dav40, Pri40, Ran42, Rou47, Sur43b, Was48]. Protected [Els41]. Protozoan [Sur40, Sur43a]. Provincial [Arm49]. Provisional [AL45]. Public [Ano49f]. Puerto [Bon41b]. Puget [PM41]. Punctatus [BH48]. Pyloric [Bel40b]. Pyramid [Sum40].

Qualifications [Dea41]. Qualitative [Nee40]. Quantitative [Nee40, Sur41]. Quebec [Ric42].

Raceways [Was48]. Rafinesque [Bec41, Bec43, Hil42, MH45]. Rainbow [GS49b, Hau49, Kla41, Law41, LL49, Mot40, Mot41, Mot42, NS45, Raw48, Ray42, TRW47, Woo43]. Raised [Har41a]. Rate [Arm49, Bec41, Bec43, Bec49, LM41, Mar41]. Rates [DF41, Kru49b, Ric49, Vio43]. Ratio [Bec49, Sur49]. Re [Sla49]. Re-formation [Sla49]. Reared [BTH48, SK48, She47, Woo43]. Rearing [Els41, Hog49a, Mee43, SM45]. Recently [RL42]. Recirculation [FB40]. Reclamation [Slo49]. Recognition [Sla49]. Recording [Sha47]. Records [Cur40]. Recovery [Bal48a]. Recreational [Pic49]. Reduction [Bec41, Bec43]. Reference [AL45, Car45a, Car45b, CGS49, O’D47, RS45, Sto47, Van48]. Refrigerated [HS41]. Regeneration [Arm49, Hea43]. Regius [Ran42]. Reighard [Job43]. reighardi [Job43]. Relation [Ano49f, CW49, HS41, Leo42, NSH41, Rus41, SL49, Smi40b, Tri41, Web48]. Relations [Dea49]. Relationship [Bec48, Car45b, Kru49a, MM45, Van42]. Relative [SF49]. Remarks [Hog49b, vO49]. Removing [RG41]. Report [Hun49]. Reports [Ano43f, Ano45d, Ano49r, Ano49u, Ano49t, Ano49s, Bag49, Dea49, Hun49, Rei49, vO47]. Requirement [PT47, PB49, Phi49]. Research [Cha43, Fel41]. Reservoir [Jon41, Lym43, LD45, Mil45, Tar42]. Reservoirs [Den48, TM43]. Resources [Fel41]. Response [GS49b]. Restoration [HR41, RS45]. Resulting [MHB48]. Results [SH41, She47, Smi41b, Sur49, Tre45, WS43]. reticulatum [O’D45b]. Returns [NS42, Smi47]. Revised [Ano48a, Ano49i, Ano49j]. Riboflavin [PT47]. Richardson [Car45b, Sum48]. Rico [Bon41b]. Riding [Raw48]. River [BH48, Bel40b, BK42, Cot49, MHB48, Pic49, Raw41, Smi41b, TRW47, WS43, NSH41, Rou47, Smi49]. Rock [Bec41, Bec43, Hil42]. Root [BB43a]. Rosyside [Sim40]. Rote [BB43a]. Rote-None [BB43a]. Rotenone [Bal48b, Smi40a, Smi41a]. Rough [Tar45]. Run [Rou47]. rupestris [Bec41, Bec43, Hil42].

Sacramento [NSH41]. Saginaw [HJ41]. Saint [Ano42h]. salar [Bel40a, Bel40b]. Salmo [Bel40a, Bel40b, BK42, EH48, Mot42, PM41, SK48, Sum48, Try43]. salmoides [SS43b, Sur43a]. Salmon [Bel40a, Bel40b, Bre48, Car40, Cha40, CGS49, DF41, HR41, NSH41, PH45, Pri40, Rou47, SB40, Sla49, Smi40b, Tin42, Smi41b]. Salmonid [FB40, M’G41]. Salmonoid [Tho40]. Salvage [NSH41]. Salvelinus [DM43, ESC^{+}41, Gut48, GS49a]. San [Ano40h]. Sandhills [Kin43]. Saskatchewan [Raw45]. Savings [CH40]. Scale [Tes41]. Scales [Mot42, Sha47]. Scientific [Ano48b]. Scyphidia [Sur40, Sur43a]. Sea [MM49, She49]. Season [HD47, Jon41]. Seasonal [LM41, NS45]. Second [Bal48a]. Secretary [Hun49]. Secretary-Treasurer [Hun49]. Section [Emb40, Sch42, SCH49]. Selecting [Rou42]. Selective [Gre41]. Sept [Ano41b]. September [Ano41h, SL43]. Service [Ano40e, Ano41e, Ano42e, Ano43e, Ano45c, Ano47f, Ano48d, Ano49p, Ano49q]. Sessional [Ano49b, Ano49c, Ano49s, Bag49]. Seth [Ano49o]. Seven [Bec48, Bec49, SS45]. Seven- [SS45]. Seventh [Ano49v]. Seventieth [Ano41h]. Seventy [Ano42h, Ano49a, Ano49v, vO47]. Seventy-First [Ano42h]. Seventy-Fourth [vO47]. Seventy-Seventh [Ano49v]. Seventy-Sixth [Ano49a]. Several [Esc43, Mil45, SS42a, SM43]. Sex [Bec49]. Sexual [Car45a, Wal41]. Shallow [Ken41]. Shasta [NSH41]. Shocker [Hau49]. Showing [Ano40c, Ano41c, Ano42c, Ano43c, Ano47d]. Shrimps [AL45]. Silt [Smi40b]. Silver [SB40]. Silvery [Ran42]. Sister [Bal47, BB43a, BB43b]. Sixth [Ano49a]. Sixty [Ano40h]. Sixty-Ninth [Ano40h]. Size [Mot41, SH41]. Sized [Sch42, She47, WS43]. Skaneateles [Ray42]. Slave [Raw49]. Slow [SK48]. Small [O’D43, O’D45b, RL42, Ric48]. Smallmouth [Bro40, Har41b, Jon41, Raw45, Sur40, Sur41, Sur43b, Sur45, Sur47, Web48]. Smelt [Van47, WFJ43]. Society [Ano40h, Ano41h, Ano42h, Ano40b, Ano41b, Ano42b, Ano43b, Ano47b, Ano49o, vO47, vO48, Ano48a, Ano49i, Ano49j]. Sockeye [Bre48, Cha40, DF41]. Sodium [Phi47]. Solutions [FB40, HWD48]. Solving [Tar45]. Some [Car49, CGS49, Hau49, Leo42, Moy49, Ric49, SJH49]. Sound [PM41]. Sources [All42]. South [Du 48]. Southern [BH48]. Spat [Bon40]. Spatterdock [SS41a]. Spawning [HD47, Ray42, Smi47]. Special [Ano40f, Ano41f, Ano42f, Ano43g, Ano45e, Ano47g, Ano48e, Ano49t, Ano49w, O’D47, Rei49, RS45, Van48]. Species [SV40, Tra41]. Speckled [Bre41, Har41a]. Spinal [O’D45a]. Spiny [Mar48]. Spirit [Ros49]. Sport [Cro40, Sum48, Tar45, Tri41]. Spring [Ken41, She47, Tre45]. St [Bel40a, Kru49a]. Stages [Wal41]. Standard [Bec48, Rod47, Roy42, Sch42]. Standardization [Hil48]. Standing [Ano40g, Ano41g, Ano42g, Ano43h, Ano45f, Ano47h, Ano48f, Ano49u, Ano49x, Dea49]. State [Eic49a, Hof41]. States [Slo49, Ano48b]. Statisticaal [Mot49]. Statistics [Cro40]. Status [Nee49b]. Steelhead [GWS41, PM41, Sum48]. Stizostedion [Car45a, Ros49, SM45]. Stock [Mot41]. Stocked [Kru49b, TRW47].

Stocking [Lan42, Ric48, Sur49]. Stocks [Van49]. Stomach [LM41]. Storage [Den48]. Storage-Type [Den48]. Stream [Cha43, Cla48, Emb40, Sch42, SL43, SCH49, Spr41]. Streams [Kla41, Nee49a, RL42, SH41, She47, Sho47b, SS45, SJH49, Sur41]. Studies [Bec43, Deg45, EH48, Hol49, Kin43, PM41, Rou42]. Study [Bjo40, BB43b, DF41, Lym43, LD45, Pri40, SL43, Sho47a, Sto47, Sur41, Rei49]. Stunted [SS42b]. Submerged [SS42a]. Success [Raw48, Wri45]. Sucker [Sim40]. Sudden [Woo42]. Suggestions [Tar42]. Sulfa [Gut48]. Sulfamerazine [GS49a, GS49b]. Sulphate [Smi40a]. Summary [Bal48b]. Summer [Ken41, SS41b]. Sunfish [Ric48]. Supersaturation [Woo42]. Supply [Cha40, Hig42]. Surfaces [Sur43a]. Survey [Dea41, Smi49]. Survival [Kru49b, Nee49a, Sch45, SK48, SS43b, SS45, Web48]. Suspension [SF49]. System [Raw41, Sha47, DM43].

T.V.A. [Esc43, TM43]. Tag [Rou42]. Tagged [Tre45]. Tagging [Bal47, Bel40a, SV40]. Taxonomy [DM43]. Technique [Fis47]. Technology [Fel41]. Temperature [Law41, Web48]. Temperatures [DF41]. Tempering [Bre41]. Ten [Hia47a, SS45]. Ten-inch [SS45]. Ten-Pounder [Hia47a]. Tennessee [Lym43, LD45, Jon41, Sho47b, SM43]. Terms [Ano40e, Ano41e, Ano42e, Ano43e, Ano45c, Ano47f, Ano48d, Ano49p, Ano49q, Rod47]. Terrace [SS40b]. Terrace-water [SS40b]. Test [HWD48]. Texas [Bon41a]. their [MA45, Sur45, Tar42]. Third [Bal47, BB43a, BB43b]. tholiformis [Sur43a]. Three [Cro40, Den48, O’D43, Sur41]. Thymallus [Leo40]. Tillamook [Sum48]. Time [Bag49]. Toronto [Ano41h]. Total [Bec48, Roy42]. Toxicity [Deg45, EH48, HWD48, SF49]. Training [Dea41]. Transactions [Ano40h, Ano41h, Ano42h, Ano48a, Ano49i, Ano49j]. Transportation [Bub40]. Transporting [Has41]. Trap [Bre48]. Treasurer [Hun49]. Treatment [Gut48, GS49a, Smi41a]. Trends [Car49, Eic49a, Hil49]. Trichodiniasis [FB40]. Troughs [Try43]. Trout [Arm49, BTH48, Bre41, BK42, CGS49, Dav40, DL41, ESC^{+}41, EH48, Fis40, FC48, GWS41, Gut48, GS49a, GS49b, Har41a, Has41, Hau49, Hea43, Hey49, Hil49, Kla41, Law41, Leo42, LL49, Mot40, Mot41, Mot42, NS42, NS45, Nee49a, O’D47, PM41, Phi47, PT47, PB49, Phi49, Pré41, RL42, Raw41, Raw48, Ray42, Sch42, Sch45, SK48, SH41, SL43, She47, SCH49, Smi40b, SV40, Smi41b, SS45, Smi47, SJH49, Spr41, Sum48, TRW47, Tre45, Try43, WS43, Wol41, Woo43]. trutta [BK42, EH48]. Tullibee [Car45b, Car45b]. Turbidity [Van48]. TVA [Den48, Tar45]. Two [CS45, SS45]. Two-Way [CS45]. Tyers [WL47]. Type [Den48].

Ulcer [Wol41]. undulatus [Wal41]. Unfertilized [How42]. United [Ano48b, Slo49]. University [WL47]. Unmarked [Arm49]. upon [FB40, Phi47, Sch42, Smi48]. Upper [NS42, Smi49]. Ureters [Du 48]. Use [Bub40, CS45, HSR^{+}42, Mot42, SS42a, BB43a, Has41, Rou42]. used [HWD48]. Useful [Luc41]. Utah [Wri45]. Utilization [DF41]. v [Ros49]. Vacuum [Hog41]. Valley [SM43]. Value [Gut48, Tar45]. Variations [LM41, Sur47]. Varieties [Bel40b]. Various [DF41, GS49b, Kat48, Sur45]. Varying [Sur49]. versus [Bre41, Roy42]. View [Vio45]. Vilas [WS43]. Virginia [HS48, Men45, SF49]. Viscera [PH45]. Vitamin [Phi49, Woo43]. vitreum [Car45a, SM45, Ros49, SM45, Car45a]. Volumetric [Tin42].

Waddell [Nee40]. Wakonichi [Ric42]. Warm [Vio45]. Warm-Water [Vio45]. Washtenaw [MH45]. Wastes [Deg45]. Water [DF41, FR45, HSR^{+}42, Ric49, SK48, Sho47b, Sto47, Sur45, Vio45, Ell42, SS40b]. Waters [Ano49f, Dav40, Esc43, Har41a, She49, SM43, Vio43]. Watts [Lym43]. Waubesa [Woo42]. Way [CS45]. Weed [Eic49b]. Weighing [GJSL41]. Weight [Bec48, Car45b, Kru49a, Law41, Mee43]. Weights [Hil48]. Weirs [CS45]. West [HS48, SF49, Lin47]. Western [NS42]. Wettable [HS48]. Wheeler [Mil45, Tar42]. which [HWD48]. White [Web43]. Whitefish [But41, SV40, Van42, VH49]. Whitmore [Tra41]. Wider [Fun49]. Wild [BTH48, Dav47, Sch45]. Wildlife [Ano49f, Cot49, Kin43]. Winter [CW49, Leo42, MH45, SS41b, Van47]. Wisconsin [Hil42, MHB48, O’D43, WS43, Woo42]. without [Mar48]. Woods [Car45a, Car45b, Car49]. Wyoming [Bjo40].

Yahara [MHB48]. Year [Tra41, Ano40c, Ano41c, Ano42c, Ano43c, Ano47d]. Years [Cro40]. Yellow [Car45a, HJ41, LM41, MH45, Ros49, SM45]. Yield [Sch42]. Yolk [Kru49b]. York [RW40, RL42, Ray42]. Young [Bre48, RW40, RL42, Web43].

Zymogen [Gag43].

References

Anderson:1945:PKS

[AL45] William W. Anderson and Milton J. Lindner. A provisional key to the shrimps of the family penaeidae with especial reference to American forms. Transactions of the American Fisheries Society, 73(1):284–319, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1942:CBF

[All42] K. Radway Allen. Comparison of bottom faunas as sources of available fish food. Transactions of the American Fisheries Society, 71(1):275–283, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:CI

[Ano40a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 69(1):301–303, January 1940. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:CLA

[Ano40b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended June 27, 1939). Transactions of the American Fisheries Society, 69(1):304–307, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:LMS

[Ano40c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 69(1): 308–325, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:O

[Ano40d] Anonymous. Officers for 1938–1939. Transactions of the American Fisheries Society, 69(1):3, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:PTS

[Ano40e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 69(1):6–7, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:SCb

[Ano40f] Anonymous. Special committees. Transactions of the American Fisheries Society, 69(1):5, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:SCa

[Ano40g] Anonymous. Standing committees, 1939–1940. Transactions of the American Fisheries Society, 69(1):4, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1940:TSN

[Ano40h] Anonymous. Transactions of the Sixty-Ninth Annual Meeting of the American Fisheries Society: San Francisco, California, June 26 and 27, 1939. Transactions of the American Fisheries Society, 69(1):13–60, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:CI

[Ano41a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 70(1):489, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:CLA

[Ano41b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended sept. 6, 1940). Transactions of the American Fisheries Society, 70(1):490–493, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:LMS

[Ano41c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 70(1): 494–511, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:O

[Ano41d] Anonymous. Officers for 1939–1940. Transactions of the American Fisheries Society, 70(1):3, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:PTS

[Ano41e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 70(1):6–7, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:SCb

[Ano41f] Anonymous. Special committee. Transactions of the American Fisheries Society, 70(1):5, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:SCa

[Ano41g] Anonymous. Standing committees, 1940–1941. Transactions of the American Fisheries Society, 70(1):4, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1941:TSA

[Ano41h] Anonymous. Transactions of the Seventieth Annual Meeting of the American Fisheries Society: Toronto, Ontario, September 5 and 6, 1940. Transactions of the American Fisheries Society, 70(1): 13–51, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:CI

[Ano42a] Anonymous. Certification of incorporation. Transactions of the American Fisheries Society, 71(1):371–373, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:CLA

[Ano42b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended August 26, 1942). Transactions of the American Fisheries Society, 71(1):374–377, January 1942. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:LMS

[Ano42c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 71(1): 378–394, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:O

[Ano42d] Anonymous. Officers for 1940–1941. Transactions of the American Fisheries Society, 71(1):3, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:PTS

[Ano42e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 71(1):6–7, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:SCb

[Ano42f] Anonymous. Special committee. Transactions of the American Fisheries Society, 71(1):5, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:SCa

[Ano42g] Anonymous. Standing committees, 1941–1942. Transactions of the American Fisheries Society, 71(1):4, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1942:TSF

[Ano42h] Anonymous. Transactions of the Seventy-First Annual Meeting of the American Fisheries Society: Saint Louis, Missouri, August 25 and 26, 1941. Transactions of the American Fisheries Society, 71(1):13–58, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:CI

[Ano43a] Anonymous. Certification of incorporation. Transactions of the American Fisheries Society, 72(1):287, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:CLA

[Ano43b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended August 26, 1942). Transactions of the American Fisheries Society, 72(1):288–291, January 1943. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:LMS

[Ano43c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 72(1): 292–308, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:O

[Ano43d] Anonymous. Officers for 1942–1943. Transactions of the American Fisheries Society, 72(1):3, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:PTS

[Ano43e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 72(1):6–7, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:RO

[Ano43f] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 72(1):11–18, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:SCb

[Ano43g] Anonymous. Special committee. Transactions of the American Fisheries Society, 72(1):5, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1943:SCa

[Ano43h] Anonymous. Standing committees, 1941–1942. Transactions of the American Fisheries Society, 72(1):4, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1945:M

[Ano45a] Anonymous. In memoriam. Transactions of the American Fisheries Society, 73(1):8, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1945:O

[Ano45b] Anonymous. Officers for 1943–1944. Transactions of the American Fisheries Society, 73(1):3, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1945:PTS

[Ano45c] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 73(1):6–7, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1945:RO

[Ano45d] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 73(1):11–15, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1945:SCb

[Ano45e] Anonymous. Special committee. Transactions of the American Fisheries Society, 73(1):5, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1945:SCa

[Ano45f] Anonymous. Standing committees, 1941–1942. Transactions of the American Fisheries Society, 73(1):4, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:CI

[Ano47a] Anonymous. Certification of incorporation. Transactions of the American Fisheries Society, 74(1):373, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:CLA

[Ano47b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended August 26, 1941). Transactions of the American Fisheries Society, 74(1):374–377, January 1947. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:M

[Ano47c] Anonymous. In memoriam. Transactions of the American Fisheries Society, 74(1):8, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:LMS

[Ano47d] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 74(1): 378–391, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:O

[Ano47e] Anonymous. Officers for 1944–1945. Transactions of the American Fisheries Society, 74(1):3, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:PTS

[Ano47f] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 74(1):6–7, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:SCb

[Ano47g] Anonymous. Special committee. Transactions of the American Fisheries Society, 74(1):5, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1947:SCa

[Ano47h] Anonymous. Standing committees, 1944–1945. Transactions of the American Fisheries Society, 74(1):4, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1948:IPM

[Ano48a] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1946). Transactions of the American Fisheries Society, 75(1):399–402, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1948:LCS

[Ano48b] Anonymous. A list of common and scientific names of the better known fishes of the United States and Canada. Transactions of the American Fisheries Society, 75(1):353–397, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1948:O

[Ano48c] Anonymous. Officers for 1945–1946. Transactions of the American Fisheries Society, 75(1):3, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1948:PTS

[Ano48d] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 75(1):6–7, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1948:SCb

[Ano48e] Anonymous. Special committee. Transactions of the American Fisheries Society, 75(1):5, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1948:SCa

[Ano48f] Anonymous. Standing committees, 1945–1946. Transactions of the American Fisheries Society, 75(1):4, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:ASS

[Ano49a] Anonymous. Agenda of Seventy-Sixth Annual Meeting. Transactions of the American Fisheries Society, 76(1):391, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:ASCa

[Ano49b] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 76(1):394, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:ASCb

[Ano49c] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 77(1):286, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:CBa

[Ano49d] Anonymous. Concluding business. Transactions of the American Fisheries Society, 76(1):443–444, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:CBb

[Ano49e] Anonymous. Concluding business. Transactions of the American Fisheries Society, 77(1):336–338, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:FWR

[Ano49f] Anonymous. Fish and wildlife: In relation to flood-control and navigation developments and to the elimination of pollution of public waters. Transactions of the American Fisheries Society, 77(1): 251–280, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:Ma

[Ano49g] Anonymous. In memoriam. Transactions of the American Fisheries Society, 76(1):8, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:Mb

[Ano49h] Anonymous. In memoriam. Transactions of the American Fisheries Society, 77(1):8, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:IPMa

[Ano49i] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1946). Transactions of the American Fisheries Society, 76(1):445–450, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:IPMb

[Ano49j] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1949). Transactions of the American Fisheries Society, 77(1):339–344, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:LMa

[Ano49k] Anonymous. List of members. Transactions of the American Fisheries Society, 76(1):451–466, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:LMb

[Ano49l] Anonymous. List of members. Transactions of the American Fisheries Society, 77(1):345–363, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:O

[Ano49m] Anonymous. Officers 1946–1947. Transactions of the American Fisheries Society, 77(1):3, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:OC

[Ano49n] Anonymous. Officers and committees, 1945–1946. Transactions of the American Fisheries Society, 76(1):3–5, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:PMS

[Ano49o] Anonymous. Presentation by Mr. Seth Gordon of Photograph of Members in Attendance at 1915 Meeting of the Society. Transactions of the American Fisheries Society, 76(1):401–403, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:PTSa

[Ano49p] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 76(1):6–7, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:PTSb

[Ano49q] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 77(1):6–7, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:RO

[Ano49r] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 76(1):395–400, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:RSCc

[Ano49s] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 76(1):433–442, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:RSCb

[Ano49t] Anonymous. Reports of special committees. Transactions of the American Fisheries Society, 76(1):425–432, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:RSCa

[Ano49u] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 76(1):404–424, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:SSA

[Ano49v] Anonymous. Seventy-Seventh Annual Meeting. Transactions of the American Fisheries Society, 77(1):281–283, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:SCb

[Ano49w] Anonymous. Special committees. Transactions of the American Fisheries Society, 77(1):5, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1949:SCa

[Ano49x] Anonymous. Standing committees, 1947–1948. Transactions of the American Fisheries Society, 77(1):4, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1949:MRG

[Arm49] George C. Armstrong. Mortality, rate of growth, and fin regeneration of marked and unmarked lake trout fingerlings at the Provincial Fish Hatchery, Port Arthur, Ontario. Transactions of the American Fisheries Society, 77(1):129–131, January 1949. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bagley:1949:RSC

[Bag49] Lester Bagley. Reports of sessional committees: Committee on Time and Place. Transactions of the American Fisheries Society, 77(1):327–335, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ball:1947:TEF

[Bal47] Robert C. Ball. A tagging experiment on the fish population of Third Sister Lake, Michigan. Transactions of the American Fisheries Society, 74(1):360–369, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ball:1948:RMF

[Bal48a] Robert C. Ball. Recovery of marked fish following a second poisoning of the population in Ford lake, Michigan. Transactions of the American Fisheries Society, 75(1):36–42, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ball:1948:SEM

[Bal48b] Robert C. Ball. A summary of experiments in Michigan lakes on the elimination of fish populations with rotenone, 1934–1942. Transactions of the American Fisheries Society, 75(1):139–146, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangham:1941:PFA

[Ban41] Ralph V. Bangham. Parasites of fish of Algonquin Park lakes. Transactions of the American Fisheries Society, 70(1):161–171, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1943:EUD

[BB43a] C. J. D. Brown and Robert C. Ball. An experiment in the use of Derris root (rote-none) on the fish and fish-food organisms of Third Sister Lake. Transactions of the American Fisheries Society, 72(1):267–284, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1943:FPS

[BB43b] C. J. D. Brown and Robert C. Ball. A fish population study of Third Sister Lake. Transactions of the American Fisheries Society, 72(1):177–186, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1941:IGR

[Bec41] William C. Beckman. Increased growth rate of rock bass, Ambloplites rupestris (Rafinesque), following reduction in the density of the population. Transactions of the American Fisheries Society, 70(1):143–148, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1943:FSI

[Bec43] William C. Beckman. Further studies on the increased growth rate of the rock bass Ambloplites rupestris (Rafinesque), following the reduction in density of the population. Transactions of the American Fisheries Society, 72(1):72–78, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1948:LWR

[Bec48] William C. Beckman. The length–weight relationship, factors for conversions between standard and total lengths, and coefficients of condition for seven Michigan fishes. Transactions of the American Fisheries Society, 75(1):237–256, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1949:RGS

[Bec49] William C. Beckman. The rate of growth and sex ratio for seven Michigan fishes. Transactions of the American Fisheries Society, 76(1):63–81, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1940:MAS

[Bel40a] David L. Belding. Migration of the Atlantic salmon (Salmo salar) in the Gulf of St. Lawrence as determined by tagging experiments. Transactions of the American Fisheries Society, 69(1):290–295, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1940:NEP

[Bel40b] David L. Belding. The number of eggs and pyloric appendages as criteria of river varieties of the Atlantic salmon (Salmo salar). Transactions of the American Fisheries Society, 69(1):285–289, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bailey:1948:FHS

[BH48] Reeve M. Bailey and Harry M. Harrison, Jr. Food habits of the southern channel catfish (Ictalurus lacustris Punctatus) in the Des Moines River, Iowa. Transactions of the American Fisheries Society, 75(1):110–138, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bjorn:1940:POE

[Bjo40] Eugene E. Bjorn. Preliminary observations and experimental study of the ling, Lota maculosa (LeSueur), in Wyoming. Transactions of the American Fisheries Society, 69(1):192–196, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1942:GME

[BK42] C. J. D. Brown and Gertrude C. Kamp. Gonad measurements and egg counts of brown trout (Salmo trutta) from the Madison River, Montana. Transactions of the American Fisheries Society, 71(1): 195–200, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonnot:1940:MCO

[Bon40] Paul Bonnot. Methods of collecting oyster spat. Transactions of the American Fisheries Society, 69(1):263–267, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonham:1941:FGT

[Bon41a] Kelshaw Bonham. Food of gars in Texas. Transactions of the American Fisheries Society, 70(1):356–362, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonnet:1941:PPC

[Bon41b] Luis C. Bonnet. Possibilities for the promotion of the commercial fisheries of Puerto Rico. Transactions of the American Fisheries Society, 70(1):95–97, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brett:1941:TVA

[Bre41] J. R. Brett. Tempering versus acclimation in the planting of speckled trout. Transactions of the American Fisheries Society, 70(1): 397–403, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brett:1948:DOT

[Bre48] J. R. Brett. The design and operation of a trap for the capture of migrating young sockeye salmon. Transactions of the American Fisheries Society, 75(1):97–104, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1940:SBB

[Bro40] Merrill W. Brown. Smallmouth black bass propagation in California. Transactions of the American Fisheries Society, 69(1):119– 124, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baeder:1948:CPH

[BTH48] Helen A. Baeder, Peter I. Tack, and Albert S. Hazzard. A comparison of the palatability of hatchery-reared and wild brook trout. Transactions of the American Fisheries Society, 75(1):181–185, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bubghduff:1940:ADU

[Bub40] A. E. Bubghduff. An aerating device for use in the transportation of fish. Transactions of the American Fisheries Society, 69(1):296– 299, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1941:AFW

[But41] George E. Butler. Artificial feeding of whitefish fry. Transactions of the American Fisheries Society, 70(1):180–182, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carbine:1948:FPD

[CA48] W. F. Carbine and Vernon C. Applegate. The fish population of Deep Lake, Michigan. Transactions of the American Fisheries Society, 75(1):200–227, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carl:1940:CCS

[Car40] G. Clifford Carl. Comparison of Coho salmon fry from eggs incubated in gravel and in hatchery baskets. Transactions of the American Fisheries Society, 69(1):132–134, January 1940. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carbine:1942:OLH

[Car42] W. F. Carbine. Observations on the life history of the northern pike, Esox lucius L., in Houghton Lake, Michigan. Transactions of the American Fisheries Society, 71(1):149–164, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1945:AGS

[Car45a] Kenneth D. Carlander. Age, growth, sexual maturity, and population fluctuations of the yellow pike-perch, Stizostedion vitreum vitreum (Mitchill), with reference to the commercial fisheries, Lake of the Woods, Minnesota. Transactions of the American Fisheries Society, 73(1):90–107, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1945:GLW

[Car45b] Kenneth D. Carlander. Growth, length–weight relationship and population fluctuations of the tullibee, Leucichthys artedi Tullibee (Richardson), with reference to the commercial fisheries, Lake of the Woods, Minnesota. Transactions of the American Fisheries Society, 73(1):125–136, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1949:STC

[Car49] Kenneth D. Carlander. Some trends in the commercial fisheries of Lake of the Woods, Minnesota. Transactions of the American Fisheries Society, 77(1):13–25, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cope:1949:ESI

[CGS49] Oliver B. Cope, Claude M. Gjullin, and Alf Storm. Effects of some insecticides on trout and salmon in Alaska, with reference to blackfly control. Transactions of the American Fisheries Society, 77(1):160–177, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1940:SGC

[CH40] G. H. Clark and S. Ross Hatton. Savings gear in the California drag-net fishery. Transactions of the American Fisheries Society, 69(1):106–110, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1940:EDO

[Cha40] Wilbert McLeod Chapman. Effects of a decreased oxygen supply on sockeye and Chinook salmon. Transactions of the American Fisheries Society, 69(1):197–204, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chamberlain:1943:RSM

[Cha43] Thomas K. Chamberlain. Research in stream management in the Pisgah National Forest. Transactions of the American Fisheries Society, 72(1):150–176, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1948:SIM

[Cla48] O. H. Clark. Stream improvements in Michigan. Transactions of the American Fisheries Society, 75(1):270–280, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cottam:1949:MRB

[Cot49] Clarence Cottam. Missouri River Basin development and its effects on fish and wildlife. Transactions of the American Fisheries Society, 76(1):380–388, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Croker:1940:TYF

[Cro40] Richard S. Croker. Three years of fisheries statistics on marine sport fishing in California. Transactions of the American Fisheries Society, 69(1):111–118, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carbine:1945:EUT

[CS45] W. F. Carbine and David S. Shetter. Examples of the use of twoway fish weirs in Michigan. Transactions of the American Fisheries Society, 73(1):70–89, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Curtis:1940:ACR

[Cur40] Brian Curtis. Anglers’ catch records in California. Transactions of the American Fisheries Society, 69(1):125–131, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1949:RDO

[CW49] Gerald P. Cooper and George N. Washburn. Relation of dissolved oxygen to winter mortality of fish in Michigan lakes. Transactions of the American Fisheries Society, 76(1):23–33, January 1949. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1940:APM

[Dav40] H. S. Davis. Artificial propagation and the management of trout waters. Transactions of the American Fisheries Society, 69(1):158– 168, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1947:IDA

[Dav47] H. S. Davis. Infectious diseases among wild fishes. Transactions of the American Fisheries Society, 74(1):21–25, January 1947. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1949:CCC

[Dav49] H. S. Davis. Cytophaga columnaris as a cause of fish epidemics. Transactions of the American Fisheries Society, 77(1):102–104, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deason:1941:SAQ

[Dea41] H. J. Deason. A survey of academic qualifications for fishery biologists and of institutional facilities for training fishery biologists. Transactions of the American Fisheries Society, 70(1):128–142, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deason:1949:RSC

[Dea49] H. J. Deason. Reports of standing committees: Committee on International Relations. Transactions of the American Fisheries Society, 77(1):304–316, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Degani:1945:STA

[Deg45] John George Degani. Studies of the toxicity of ammunition plant wastes to fishes. Transactions of the American Fisheries Society, 73(1):45–51, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dendy:1948:PDD

[Den48] Jack S. Dendy. Predicting depth distribution of fish in three TVA storage-type reservoirs. Transactions of the American Fisheries Society, 75(1):65–71, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1941:ESE

[DF41] Lauren R. Donaldson and Fred J. Foster. Experimental study of the effect of various water temperatures on the growth, food utilization, and mortality rates of fingerling sockeye salmon. Transactions of the American Fisheries Society, 70(1):339–346, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deason:1947:AGK

[DH47] Hilary J. Deason and Ralph Hile. Age and growth of the kiyi, Leucichthys kiyi Koelz, in Lake Michigan. Transactions of the American Fisheries Society, 74(1):88–142, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1941:NFD

[DL41] H. S. Davis and Estelle C. Lazar. A new fungus disease of trout. Transactions of the American Fisheries Society, 70(1):264–271, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeLacy:1943:THC

[DM43] Allan C. DeLacy and W. Markham Morton. Taxonomy and habits of the charrs, Salvelinus malma and Salvelinus alpinus, of the Karluk Drainage System. Transactions of the American Fisheries Society, 72(1):79–91, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DuPlessis:1948:GPU

[Du 48] S. S. Du Plessis. A gyrodactyloid parasite from the ureters of largemouth bass at the Jonkershoek Inland Fish Hatchery, South Africa. Transactions of the American Fisheries Society, 75(1):105– 109, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Everhart:1948:AST

[EH48] W. Harry Everhart and William W. Hassler. Aquarium studies on the toxicity of DDT to brown trout, Salmo trutta. Transactions of the American Fisheries Society, 75(1):59–64, January 1948. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eicher:1949:CTS

[Eic49a] George J. Eicher, Jr. Current trends in State fishery programs. Transactions of the American Fisheries Society, 76(1):13–22, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eicher:1949:LWC

[Eic49b] George J. Eicher, Jr. Localized weed control in management of game fish. Transactions of the American Fisheries Society, 76(1): 177–182, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1942:FWI

[Ell42] M. M. Ellis. Fresh-water impoundments. Transactions of the American Fisheries Society, 71(1):80–93, January 1942. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elson:1941:RMP

[Els41] Paul F. Elson. Rearing maskinonge in a protected area. Transactions of the American Fisheries Society, 70(1):421–429, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1940:MEN

[Emb40] Daniel R. Embody. A method of estimating the number of fish in a given section of a stream. Transactions of the American Fisheries Society, 69(1):231–236, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1942:MMN

[Emb42] Daniel R. Embody. A method of measuring the number of ectoparasites infesting hatchery fish, with particular application to Gyrodactylus. Transactions of the American Fisheries Society, 71 (1):122–130, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1941:ECB

[ESC^{+}41] Daniel R. Embody, H. A. Schuck, S. Lee Crump, J. W. Freese, and Lionel Ross. The effect of cresol on brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 70(1): 304–310, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1943:EIF

[Esc43] R. W. Eschmeyer. The effect of impoundment on fishing intensity in several T.V.A. waters. Transactions of the American Fisheries Society, 72(1):103–107, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1942:AMP

[EWS^{+}42] Chairman R. W. Eschmeyer, Herbert E. Warfel, Lloyd L. Smith, Jr., C. J. D. Brown, Stillman Wright, A. Randolph Shields, Lyle M. Thorpe, and H. S. Swingle. An appraisal of management policies for lakes and ponds. Transactions of the American Fisheries Society, 71(1):332–342, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1940:EUC

[FB40] Frederic F. Fish and Roger E. Burrows. Experiments upon the control of trichodiniasis of salmonid fishes by the prolonged recirculation of formalin solutions. Transactions of the American Fisheries Society, 69(1):94–100, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fry:1948:LTF

[FC48] F. E. J. Fry and V. B. Chapman. The lake trout fishery in Algonquin Park from 1936 to 1945. Transactions of the American Fisheries Society, 75(1):19–35, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fellers:1941:RFT

[Fel41] Carl R. Fellers. Research in food technology in the development of our fishery resources. Transactions of the American Fisheries Society, 70(1):72–76, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1940:ETC

[Fis40] Frederic F. Fish. An evaluation of trout culture. Transactions of the American Fisheries Society, 69(1):85–89, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1941:NCN

[Fis41] Frederic F. Fish. Notes on Costia necatrix. Transactions of the American Fisheries Society, 70(1):441–445, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1943:AFH

[Fis43] Frederic F. Fish. The anaesthesia of fish by high carbon-dioxide concentrations. Transactions of the American Fisheries Society, 72(1):25–29, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1947:TCI

[Fis47] Frederic F. Fish. A technique for controlling infectious disease in hatchery fish. Transactions of the American Fisheries Society, 74 (1):209–222, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1945:CDC

[FR45] Frederic F. Fish and Robert R. Rucker. Columnaris as a disease of cold-water fishes. Transactions of the American Fisheries Society, 73(1):32–36, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Funk:1949:WAE

[Fun49] John L. Funk. Wider application of the electrical method of collecting fish. Transactions of the American Fisheries Society, 77(1): 49–60, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gage:1943:ZGF

[Gag43] Simon Henry Gage. Zymogen granules in the fishes. Transactions of the American Fisheries Society, 72(1):263–266, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffiths:1941:MFL

[GJSL41] Francis P. Griffiths, Willard Jarvis, Andrew B. Smith, and Charles A. Lockwood. Measurement of fish for liberation by weighing. Transactions of the American Fisheries Society, 70(1):275– 281, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greenbank:1941:SPF

[Gre41] John Greenbank. Selective poisoning of fish. Transactions of the American Fisheries Society, 70(1):80–86, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffiths:1940:CID

[Gri40] Francis P. Griffiths. Considerations of the introduction and distribution of exotic fishes in Oregon. Transactions of the American Fisheries Society, 69(1):240–243, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gutsell:1949:DST

[GS49a] James S. Gutsell and S. F. Snieszko. Dosage of sulfamerazine in the treatment of furunculosis in brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 76(1):82–96, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gutsell:1949:RBR

[GS49b] James S. Gutsell and S. F. Snieszko. Response of brook, rainbow, and brown trout to various dosages of sulfamerazine. Transactions of the American Fisheries Society, 77(1):93–101, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gutsell:1948:VCD

[Gut48] James S. Gutsell. The value of certain drugs, especially sulfa drugs, in the treatment of furunculosis in brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 75(1):186– 199, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffiths:1941:EAS

[GWS41] Francis P. Griffiths, Glaucus Webb, and Phillip W. Schneider. Ether anesthesia of steelhead trout. Transactions of the American Fisheries Society, 70(1):272–274, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harkness:1941:CST

[Har41a] William J. K. Harkness. Catches of speckled trout from the plantings of hatchery-raised fish in private waters of Ontario. Transactions of the American Fisheries Society, 70(1):410–413, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hart:1941:ASB

[Har41b] J. S. Hart. The availability of smallmouth black bass, Micropterus dolomieu Lacépède, in lakes of Algonquin Park, Ontario. Transactions of the American Fisheries Society, 70(1):172–179, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haskell:1940:EMC

[Has40] David C. Haskell. An electrical method of collecting fish. Transactions of the American Fisheries Society, 69(1):210–215, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haskell:1941:IUO

[Has41] David C. Haskell. An investigation of the use of oxygen in transporting trout. Transactions of the American Fisheries Society, 70 (1):149–160, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hauck:1949:SHE

[Hau49] Forrest R. Hauck. Some harmful effects of the electric shocker on large rainbow trout. Transactions of the American Fisheries Society, 77(1):61–64, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1947:DAS

[HD47] Ralph Hile and Hilary J. Deason. Distribution, abundance, and spawning season and grounds of the kiyi, Leucichthys kiyi Koelz, in Lake Michigan. Transactions of the American Fisheries Society, 74(1):143–165, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heacox:1943:FRB

[Hea43] Cecil Heacox. Fin regeneration in brown trout. Transactions of the American Fisheries Society, 72(1):231–232, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hey:1949:FBT

[Hey49] D. Hey. The fertility of brown trout eggs at the jonkershoek inland fish hatchery. Transactions of the American Fisheries Society, 77(1):65–80, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hiatt:1947:FCFa

[Hia47a] Robert W. Hiatt. Food-chains and the food cycle in Hawaiian fish ponds. — Part I. The food and feeding habits of mullet (Mugil cephalus), milkfish (Chanos chanos), and the ten-pounder (Elops machnata). Transactions of the American Fisheries Society, 74(1): 250–261, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hiatt:1947:FCFb

[Hia47b] Robert W. Hiatt. Food-chains and the food cycle in Hawaiian fish ponds. — Part II. Biotic interaction. Transactions of the American Fisheries Society, 74(1):262–280, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Higgins:1942:CFS

[Hig42] Elmer Higgins. Can the fisheries supply more food during a national emergency? Transactions of the American Fisheries Society, 71(1):61–73, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1942:GRB

[Hil42] Ralph Hile. Growth of the rock bass, Ambloplites rupestris (Rafinesque), in five lakes of Northeastern Wisconsin. Transactions of the American Fisheries Society, 71(1):131–143, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1948:SME

[Hil48] Ralph Hile. Standardization of methods of expressing lengths and weights of fish. Transactions of the American Fisheries Society, 75(1):157–164, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1949:TLT

[Hil49] Ralph Hile. Trends in the lake trout fishery of Lake Huron through 1946. Transactions of the American Fisheries Society, 76(1):121– 147, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1941:AGP

[HJ41] Ralph Hile and Frank W. Jobes. Age, growth, and production of the yellow perch, Perca flavescens (Mitchill), of Saginaw Bay. Transactions of the American Fisheries Society, 70(1):102–122, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffmann:1941:IFS

[Hof41] Henry J. Hoffmann. Inspection of fish by the State of Minnesota. Transactions of the American Fisheries Society, 70(1):483–485, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogan:1941:EHV

[Hog41] Joe Hogan. The effects of high vacuum on fish. Transactions of the American Fisheries Society, 70(1):469–474, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogan:1949:CAP

[Hog49a] Joe Hogan. The control of aquatic plants with fertilizers in rearing ponds at the lonoke hatchery, Arkansas. Transactions of the American Fisheries Society, 76(1):183–189, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogan:1949:ORP

[Hog49b] Joe Hogan. Opening remarks of the President. Transactions of the American Fisheries Society, 77(1):284–285, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holloway:1949:LMS

[Hol49] R. C. Holloway. Lake-management studies in Oregon. Transactions of the American Fisheries Society, 77(1):105–113, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howell:1942:BOF

[How42] Henry H. Howell. Bottom organisms in fertilized and unfertilized fish ponds in Alabama. Transactions of the American Fisheries Society, 71(1):165–179, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunter:1940:PNP

[HR40] George W. Hunter III and J. Stewart Rankin, Jr. Parasites of northern pike and pickerel. Transactions of the American Fisheries Society, 69(1):268–272, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herrington:1941:RAS

[HR41] William C. Herrington and George A. Rounsefell. Restoration of the Atlantic salmon in New England. Transactions of the American Fisheries Society, 70(1):123–127, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrison:1941:RRL

[HS41] Roger W. Harrison and Maurice E. Stansby. The relation of refrigerated lockers to conservation of fishery products. Transactions of the American Fisheries Society, 70(1):260–263, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffmann:1948:EAA

[HS48] Clarence H. Hoffmann and Eugene W. Surber. Effects of an aerial application of wettable DDT on fish and fish-food organisms in Back Creek, West Virginia. Transactions of the American Fisheries Society, 75(1):48–58, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1942:PFP

[HSR^{+}42] Chairman Carl L. Hubbs, H. L. Shantz, Kenneth A. Reed, J. N. Darling, Frank B. Wire, A. H. Wiebe, James R. Simon, Elmer Higgins, and A. D. Aldrich. The place of fish production in a program of multiple water use. Transactions of the American Fisheries Society, 71(1):297–314, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hammer:1942:CFI

[HT42] Ralph C. Hammer and R. V. Truitt. Control of fishing intensity in Maryland. Transactions of the American Fisheries Society, 71(1): 144–148, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1948:FF

[Hun48] A. G. Huntsman. Freshets and fish. Transactions of the American Fisheries Society, 75(1):257–266, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunter:1949:ROR

[Hun49] Russell P. Hunter. Reports of officers: Report of the Secretary-Treasurer. Transactions of the American Fisheries Society, 77(1): 287–303, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hart:1948:AOT

[HWD48] W. B. Hart, Roy F. Weston, and J. G. Demann. An apparatus for oxygenating test solutions in which fish are used as test animals for evaluating toxicity. Transactions of the American Fisheries Society, 75(1):228–236, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haskell:1941:FDE

[HZ41] David C. Haskell and Robert G. Zilliox. Further developments of the electrical method of collecting fish. Transactions of the American Fisheries Society, 70(1):404–409, January 1941. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1943:GCL

[Jam43] M. C. James. Glen c. Leach — 1872–1942. Transactions of the American Fisheries Society, 72(1):19–24, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jobes:1943:AGB

[Job43] Frank W. Jobes. The age, growth, and bathymetric distribution of Reighard’s chub, Leucichthys reighardi Koelz, in Lake Michigan. Transactions of the American Fisheries Society, 72(1):108– 135, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jobes:1949:AGD

[Job49] Frank W. Jobes. The age, growth, and distribution of the longjaw cisco, Leucichthys alpenae Koelz, in Lake Michigan. Transactions of the American Fisheries Society, 76(1):215–247, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1941:LGS

[Jon41] Alden M. Jones. The length of the growing season of largemouth and smallmouth black bass in Norris Reservoir, Tennessee. Transactions of the American Fisheries Society, 70(1):183–187, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1948:FHV

[Kat48] Max Katz. The fecundity of herring from various parts of the North Pacific. Transactions of the American Fisheries Society, 75(1):72– 76, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kennedy:1941:MFS

[Ken41] W. A. Kennedy. The migration of fish from a shallow to a deep lake in spring and early summer. Transactions of the American Fisheries Society, 70(1):391–396, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:1943:LMS

[Kin43] Willis King. Lake management studies in the Sandhills Wildlife Management Area. Transactions of the American Fisheries Society, 72(1):204–211, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klak:1940:NIB

[Kla40] George E. Klak. Neascus infestation of black-head, blunt-nosed, and other forage minnows. Transactions of the American Fisheries Society, 69(1):273–278, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klak:1941:CBT

[Kla41] George E. Klak. The condition of brook trout and rainbow trout from four Eastern streams. Transactions of the American Fisheries Society, 70(1):282–289, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krumholz:1949:LWR

[Kru49a] Louis A. Krumholz. Length–weight relationship of the muskellunge, Esox m. Masquinongy, in Lake St. Clair. Transactions of the American Fisheries Society, 77(1):42–48, January 1949. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krumholz:1949:RSG

[Kru49b] Louis A. Krumholz. Rates of survival and growth of bluegill yolk fry stocked at different intensities in hatchery ponds. Transactions of the American Fisheries Society, 76(1):190–203, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1942:FSP

[Lan42] Chairman T. H. Langlois. Fish-stocking policies and programs. Transactions of the American Fisheries Society, 71(1):343–370, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lawrence:1941:ETW

[Law41] W. M. Lawrence. The effect of temperature on the weight of fasting rainbow trout fingerlings. Transactions of the American Fisheries Society, 70(1):290–296, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lyman:1945:PIB

[LD45] F. Earle Lyman and Jack S. Dendy. A pre-impoundment bottomfauna study of Cherokee Reservoir Area (Tennessee). Transactions of the American Fisheries Society, 73(1):194–208, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leonard:1940:FOF

[Leo40] Justin W. Leonard. Further observations on the feeding habits of the Montana grayling (Thymallus montanus) and the bluegill (Lepomis macrochirus) in Ford Lake, Michigan. Transactions of the American Fisheries Society, 69(1):244–256, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leonard:1942:SOW

[Leo42] Justin W. Leonard. Some observations on the winter feeding habits of brook trout fingerlings in relation to natural food organisms present. Transactions of the American Fisheries Society, 71(1): 219–227, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lindner:1947:CMF

[Lin47] Milton J. Lindner. The commercial marine fishes, crustaceans and molluscs of the West Coast of Mexico. Transactions of the American Fisheries Society, 74(1):71–80, January 1947. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leonard:1949:AFH

[LL49] J. W. Leonard and F. A. Leonard. An analysis of the feeding habits of rainbow trout and lake trout in Birch Lake, Cass County, Michigan. Transactions of the American Fisheries Society, 76(1): 301–314, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langford:1941:SVS

[LM41] R. R. Langford and W. R. Martin. Seasonal variations in stomach contents and rate of growth in a population of yellow perch. Transactions of the American Fisheries Society, 70(1):436–440, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lucas:1941:HFM

[Luc41] Clarence Remington Lucas. How fishery market news information is useful in conservation. Transactions of the American Fisheries Society, 70(1):87–94, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lyman:1943:PIB

[Lym43] F. Earle Lyman. A pre-impoundment bottom-fauna study of Watts Bar Reservoir Area (Tennessee). Transactions of the American Fisheries Society, 72(1):52–62, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1945:IFN

[MA45] Robert R. Miller and J. R. Alcorn. The introduced fishes of Nevada, with a history of their introduction. Transactions of the American Fisheries Society, 73(1):173–193, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1941:RGL

[Mar41] William R. Martin. Rate of growth of the ling, Lota lota Maculosa (LeSueur). Transactions of the American Fisheries Society, 70(1): 77–79, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marshall:1948:MGS

[Mar48] Nelson Marshall. The molting without growth of spiny lobsters, Panulirus argus, kept in a live car. Transactions of the American Fisheries Society, 75(1):267–269, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehean:1942:FPF

[Mee42] O. Lloyd Meehean. Fish populations of five Florida lakes. Transactions of the American Fisheries Society, 71(1):184–194, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehean:1943:GWP

[Mee43] O. Lloyd Meehean. Gain in weight per day as a measure of production in fish rearing ponds. Transactions of the American Fisheries Society, 72(1):220–230, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Menzel:1945:CFV

[Men45] R. Winston Menzel. The catfish fishery of Virginia. Transactions of the American Fisheries Society, 73(1):364–372, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MGonigle:1941:ACE

[M’G41] R. H. M’Gonigle. Acute catarrhal enteritis of salmonid fingerlings. Transactions of the American Fisheries Society, 70(1):297– 303, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moffett:1945:WFH

[MH45] James W. Moffett and Burton P. Hunt. Winter feeding habits of bluegills, Lepomis macrochirus Rafinesque, and yellow perch, Perca flavescens (Mitchill), in Cedar Lake, Washtenaw County, Michigan. Transactions of the American Fisheries Society, 73(1): 231–242, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mackenthun:1948:HMF

[MHB48] Kenneth M. Mackenthun, Elmer F. Herman, and Alfred F. Bartsch. A heavy mortality of fishes resulting from the decomposition of algae in the Yahara River, Wisconsin. Transactions of the American Fisheries Society, 75(1):175–180, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1945:CHN

[Mil45] Lawrence F. Miller. A comparison of the hoop-net catches in several fish habitats of Wheeler Reservoir. Transactions of the American Fisheries Society, 73(1):37–40, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehean:1945:RBP

[MM45] O. Lloyd Meehean and Francis Marzulli. The relationship between the production of fish and the carbon and nitrogen contents of fertilized fish ponds. Transactions of the American Fisheries Society, 73(1):262–273, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacKay:1949:RIS

[MM49] H. H. MacKay and Earl MacGillivray. Recent investigations on the sea lamprey, Petromyzon marinus, in Ontario. Transactions of the American Fisheries Society, 76(1):148–159, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1940:PRT

[Mot40] C. McC. Mottley. The production of rainbow trout at Paul Lake, British Columbia. Transactions of the American Fisheries Society, 69(1):187–191, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1941:EIS

[Mot41] C. McC. Mottley. The effect of increasing the stock in a lake on the size and condition of rainbow trout. Transactions of the American Fisheries Society, 70(1):414–420, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1942:USR

[Mot42] C. McC. Mottley. The use of the scales of rainbow trout (Salmo gairdnerii) to make direct comparisons of growth. Transactions of the American Fisheries Society, 71(1):74–79, January 1942. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1949:SAC

[Mot49] Charles M. Mottley. The statisticaal analysis of creel-census data. Transactions of the American Fisheries Society, 76(1):290–300, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1949:SIL

[Moy49] John B. Moyle. Some indices of lake productivity. Transactions of the American Fisheries Society, 76(1):322–334, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1940:QQO

[Nee40] P. R. Needham. Quantitative and qualitative observations on fish foods in Waddell Creek Lagoon. Transactions of the American Fisheries Society, 69(1):178–186, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1949:STS

[Nee49a] Paul R. Needham. Survival of trout in streams. Transactions of the American Fisheries Society, 77(1):26–31, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neess:1949:DSP

[Nee49b] John C. Neess. Development and status of pond fertilization in Central Europe. Transactions of the American Fisheries Society, 76(1):335–358, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1942:FMP

[NS42] Paul R. Needham and Frank K. Sumner. Fish management problems of high western lakes with returns from marked trout planted in Upper Angora Lake California. Transactions of the American Fisheries Society, 71(1):249–269, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1945:SCG

[NS45] P. R. Needham and Daniel W. Slater. Seasonal changes in growth, mortality, and condition of rainbow trout following planting. Transactions of the American Fisheries Society, 73(1):117–124, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1941:SSP

[NSH41] Paul R. Needham, Osgood R. Smith, and Harry A. Hanson. Salmon salvage problems in relation to Shasta Dam, California, and notes on the biology of the Sacramento River salmon. Transactions of the American Fisheries Society, 70(1):55–69, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

ODonnell:1943:FPT

[O’D43] D. John O’Donnell. The fish population in three small lakes in Northern Wisconsin. Transactions of the American Fisheries Society, 72(1):187–196, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

ODonnell:1945:COS

[O’D45a] D. John O’Donnell. A case of ossification of the spinal column in fishes. Transactions of the American Fisheries Society, 73(1): 41–44, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

ODonnell:1945:CHR

[O’D45b] D. John O’Donnell. Control of Hydrodictyon reticulatum in small ponds. Transactions of the American Fisheries Society, 73(1):59– 62, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

ODonnell:1947:DMT

[O’D47] D. John O’Donnell. The disinfection and maintenance of trout hatcheries for the control of disease, with special reference to furunculosis. Transactions of the American Fisheries Society, 74(1): 26–34, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1949:NBR

[PB49] Arthur M. Phillips, Jr. and Donald R. Brockway. The niacin and biotin requirement of trout. Transactions of the American Fisheries Society, 77(1):152–159, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1945:ESF

[PH45] Arthur M. Phillips, Jr. and George S. Hewitt. Evaluation of salmon flesh and salmon viscera in the diet of Chinook salmon fingerlings. Transactions of the American Fisheries Society, 73(1):63–69, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1947:PES

[Phi47] Arthur M. Phillips, Jr. The physiological effect of sodium chloride upon brook trout. Transactions of the American Fisheries Society, 74(1):297–309, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1949:VBR

[Phi49] Arthur M. Phillips, Jr. The vitamin B requirement of trout. Transactions of the American Fisheries Society, 76(1):34–45, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pick:1949:RFM

[Pic49] Brigadier-General Lewis A. Pick. Recreational features of Missouri River Development Program. Transactions of the American Fisheries Society, 76(1):361–372, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pautzke:1941:SLH

[PM41] Clarence F. Pautzke and Robert C. Meigs. Studies on the life history of the Puget Sound steelhead trout (Salmo gairdnerii). Transactions of the American Fisheries Society, 70(1):209–220, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prevost:1941:MIC

[Pré41] Gustave Prévost. A method of increasing the capacity of a trout hatchery. Transactions of the American Fisheries Society, 70(1): 430–435, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pritchard:1940:SNP

[Pri40] Andrew L. Pritchard. A study of the natural propagation of the pink salmon, Oncorhynchus gorbuscha, in British Columbia. Transactions of the American Fisheries Society, 69(1):237–239, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1947:RPA

[PT47] Arthur M. Phillips, Jr. and A. V. Tunison. The riboflavin and pantothenic acid requirement of brook trout. Transactions of the American Fisheries Society, 74(1):81–87, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raney:1942:PSM

[Ran42] Edward C. Raney. Propagation of the silvery minnow (Hybognathus nuchalis Regius Girard) in ponds. Transactions of the American Fisheries Society, 71(1):215–218, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1941:EBT

[Raw41] Donald S. Rawson. The Eastern brook trout in the Maligne River System, Jasper National Park. Transactions of the American Fisheries Society, 70(1):221–235, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1945:EIS

[Raw45] Donald S. Rawson. The experimental introduction of smallmouth black bass into lakes of the Prince Albert National Park, Saskatchewan. Transactions of the American Fisheries Society, 73(1):19–31, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1948:FRT

[Raw48] D. S. Rawson. The failure of rainbow trout and initial success with the introduction of lake trout in Clear Lake, Riding Mountain Park, Manitoba. Transactions of the American Fisheries Society, 75(1): 323–335, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1949:EFP

[Raw49] D. S. Rawson. Estimating the fish production of Great Slave Lake. Transactions of the American Fisheries Society, 77(1):81–92, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rayner:1942:SMR

[Ray42] H. J. Rayner. The spawning migration of rainbow trout at Skaneateles Lake, New York. Transactions of the American Fisheries Society, 71(1):180–183, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:1949:RSC

[Rei49] Kenneth A. Reid. Reports of special committees: Committee on Pollution Study. Transactions of the American Fisheries Society, 77(1):317–326, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ricker:1941:ERC

[RG41] William E. Ricker and John Gottschalk. An experiment in removing coarse fish from a lake. Transactions of the American Fisheries Society, 70(1):382–390, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richardson:1942:PFL

[Ric42] Laurence R. Richardson. The parasites of the fishes of Lake Wakonichi, Central Northern Quebec. Transactions of the American Fisheries Society, 71(1):286–289, January 1942. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ricker:1948:HSS

[Ric48] William E. Ricker. Hybrid sunfish for stocking small ponds. Transactions of the American Fisheries Society, 75(1):84–96, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ricker:1949:MRS

[Ric49] William E. Ricker. Mortality rates in some little-exploited populations of fresh-water fishes. Transactions of the American Fisheries Society, 77(1):114–128, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rounsefell:1945:HMF

[RK45] George A. Rounsefell and John Lawrence Kask. How to mark fish. Transactions of the American Fisheries Society, 73(1):320– 363, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raney:1942:AFR

[RL42] Edward C. Raney and Ernest A. Lachner. Autumn food of recently planted young brown trout in small streams of Central New York. Transactions of the American Fisheries Society, 71(1):106– 111, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ricker:1945:PIH

[RML45] William E. Ricker, Harrell F. Mosbaugh, and Maurice Lung. Production of Indiana hatcheries in 1942. Transactions of the American Fisheries Society, 73(1):373–376, January 1945. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodd:1947:STM

[Rod47] James A. Rodd. Standard terms and methods of measurement in fish culture. Transactions of the American Fisheries Society, 74(1): 19–20, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1949:PYP

[Ros49] Earl T. Rose. The population of yellow pikeperch (Stizostedion v. Vitreum) in Spirit Lake, Iowa. Transactions of the American Fisheries Society, 77(1):32–41, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rounsefell:1942:FES

[Rou42] George A. Rounsefell. Field experiments in selecting the most efficient tag for use in haddock studies. Transactions of the American Fisheries Society, 71(1):228–235, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roundsefell:1947:ENA

[Rou47] George A. Roundsefell. The effect of natural and artificial propagation in maintaining a run of Atlantic salmon in the Penobscot River. Transactions of the American Fisheries Society, 74(1):188– 208, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Royce:1942:SLV

[Roy42] William F. Royce. Standard length versus total length. Transactions of the American Fisheries Society, 71(1):270–274, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodd:1942:CFC

[RS42] James A. Rodd and Charles F. Stapledon. Classification of fishcultural products and methods of determining the numbers of fish contained in individual allotments. Transactions of the American Fisheries Society, 71(1):290–293, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rounsefell:1945:RMN

[RS45] George A. Rounsefell and Louis D. Stringer. Restoration and management of the New England alewife fisheries with special reference to Maine. Transactions of the American Fisheries Society, 73(1): 394–424, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Russell:1941:FCA

[Rus41] Ralph Russell. Fishery cooperative associations in relation to conservation and orderly marketing. Transactions of the American Fisheries Society, 70(1):188–192, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raney:1940:FGY

[RW40] Edward C. Raney and Dwight A. Webster. The food and growth of the young of the common bullhead, Ameiurus nebulosus nebulosus (LeSueur), in Cayuga Lake, New York. Transactions of the American Fisheries Society, 69(1):205–209, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shapovalov:1940:EHS

[SB40] Leo Shapovalov and William Berrian. An experiment in hatching silver salmon (Oncorhynchus kisutch) eggs in gravel. Transactions of the American Fisheries Society, 69(1):135–140, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schuck:1942:EPD

[Sch42] Howard A. Schuck. The effect of population density of legal-sized trout upon the yield per standard fishing effort in a controlled section of stream. Transactions of the American Fisheries Society, 71(1):236–248, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schuck:1945:SPD

[Sch45] Howard A. Schuck. Survival, population density, growth, and movement of the wild brown trout in Crystal Creek. Transactions of the American Fisheries Society, 73(1):209–230, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1949:EDS

[SCH49] David S. Shetter, O. H. Clark, and Albert S. Hazzard. The effects of deflectors in a section of a Michigan trout stream. Transactions of the American Fisheries Society, 76(1):248–278, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1949:RTS

[SF49] Eugene W. Surber and Dorothy D. Friddle. Relative toxicity of suspension and oil formulations of DDT to native fishes in Back Creek, West Virginia. Transactions of the American Fisheries Society, 76 (1):315–321, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1941:RPM

[SH41] David S. Shetter and Albert S. Hazzard. Results from plantings of marked trout of legal size in streams and lakes of Michigan. Transactions of the American Fisheries Society, 70(1):446–468, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaw:1940:PCC

[Sha40] Paul A. Shaw. Pollution control in California. Transactions of the American Fisheries Society, 69(1):90–93, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shapovalov:1947:SRM

[Sha47] Leo Shapovalov. A system for recording measurements of fish scales. Transactions of the American Fisheries Society, 74(1):59– 62, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1947:FRS

[She47] David S. Shetter. Further results from spring and fall plantings of legal-sized, hatchery-reared trout in streams and lakes of Michigan. Transactions of the American Fisheries Society, 74(1):35–58, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1949:BHS

[She49] David S. Shetter. A brief history of the sea lamprey problem in Michigan waters. Transactions of the American Fisheries Society, 76(1):160–176, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoemaker:1947:LSF

[Sho47a] Hurst H. Shoemaker. A laboratory study of fish populations. Transactions of the American Fisheries Society, 74(1):350–359, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:1947:GIW

[Sho47b] C. S. Shoup. Geochemical interpretation of water analyses from Tennessee streams. Transactions of the American Fisheries Society, 74(1):223–229, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simpson:1940:POE

[Sim40] James C. Simpson. Preliminary observations on the egg-eating habits of the rosyside sucker. Transactions of the American Fisheries Society, 69(1):176–177, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1949:FPS

[SJH49] Lloyd L. Smith, Jr., Raymond E. Johnson, and Laurence Hiner. Fish populations in some Minnesota trout streams. Transactions of the American Fisheries Society, 76(1):204–214, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schuck:1948:SGF

[SK48] Howard A. Schuck and Oliver R. Kingsbury. Survival and growth of fingerling brown trout (Salmo fario) reared under different hatchery conditions and planted in fast and slow water. Transactions of the American Fisheries Society, 75(1):147–156, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1943:PSL

[SL43] David S. Shetter and Justin W. Leonard. A population study of a limited area in a Michigan trout stream, September, 1940. Transactions of the American Fisheries Society, 72(1):35–51, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Salyer:1949:EBK

[SL49] J. Clark Salyer II and Karl F. Lagler. The eastern belted kingfisher, Megaceryle alcyon Alcyon (Linnaeus), in relation to fish management. Transactions of the American Fisheries Society, 76(1):97– 117, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Slater:1949:RFE

[Sla49] Daniel W. Slater. Re-formation of excised fins of king salmon fingerlings and its effects on recognition of marked adults. Transactions of the American Fisheries Society, 77(1):132–140, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sloan:1949:PPU

[Slo49] William Sloan. Program and policies of the United States Bureau of Reclamation. Transactions of the American Fisheries Society, 76(1):373–380, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1943:CHN

[SM43] Charles G. Smith and Lawrence F. Miller. A comparison of the hoop-net catch on several waters in the Tennessee Valley before and after impoundment. Transactions of the American Fisheries Society, 72(1):212–219, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1945:FIP

[SM45] Lloyd L. Smith, Jr. and John B. Moyle. Factors influencing production of yellow pikeperch, Stizostedion vitreum Vitreum, in Minnesota rearing ponds. Transactions of the American Fisheries Society, 73(1):243–261, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1940:CSR

[Smi40a] M. W. Smith. Copper sulphate and rotenone as fish poisons. Transactions of the American Fisheries Society, 69(1):141–157, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1940:PMS

[Smi40b] Osgood R. Smith. Placer mining silt and its relation to salmon and trout on the Pacific Coast. Transactions of the American Fisheries Society, 69(1):225–230, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1941:TPL

[Smi41a] M. W. Smith. Treatment of Potter’s Lake, New Brunswick, with rotenone. Transactions of the American Fisheries Society, 70(1): 347–355, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1941:RPB

[Smi41b] Lloyd L. Smith, Jr. The results of planting brook trout of legal length in the Salmon Trout River, Northern Michigan. Transactions of the American Fisheries Society, 70(1):249–259, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1947:RNS

[Smi47] Osgood R. Smith. Returns from natural spawning of cutthroat trout and Eastern brook trout. Transactions of the American Fisheries Society, 74(1):281–296, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1948:POU

[Smi48] M. W. Smith. Preliminary observations upon the fertilization of Crecy Lake, New Brunswick. Transactions of the American Fisheries Society, 75(1):165–174, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1949:CFS

[Smi49] Lloyd L. Smith, Jr. Cooperative fishery survey of the Upper Mississippi River. Transactions of the American Fisheries Society, 76 (1):279–282, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sprules:1941:EBD

[Spr41] William M. Sprules. The effect of a beaver dam on the insect fauna of a trout stream. Transactions of the American Fisheries Society, 70(1):236–248, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1940:EOI

[SS40a] E. V. Smith and H. S. Swingle. Effect of organic and inorganic fertilizers on plankton production and bluegill bream carrying capacity of ponds. Transactions of the American Fisheries Society, 69(1):257–262, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1940:FPT

[SS40b] H. S. Swingle and E. V. Smith. Fish production in terrace-water ponds in Alabama. Transactions of the American Fisheries Society, 69(1):101–105, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1941:CSN

[SS41a] E. V. Smith and H. S. Swingle. Control of spatterdock (Nuphar advena Ait.) in ponds. Transactions of the American Fisheries Society, 70(1):363–368, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1941:WSG

[SS41b] E. V. Smith and H. S. Swingle. Winter and summer growth of bluegills in fertilized ponds. Transactions of the American Fisheries Society, 70(1):335–338, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1942:UFC

[SS42a] E. V. Smith and H. S. Swingle. The use of fertilizer for controlling several submerged aquatic plants in ponds. Transactions of the American Fisheries Society, 71(1):94–101, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1942:MPS

[SS42b] H. S. Swingle and E. V. Smith. The management of ponds with stunted fish populations. Transactions of the American Fisheries Society, 71(1):102–105, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1943:OMF

[SS43a] E. V. Smith and H. S. Swingle. Organic materials as fertilizers for fish ponds. Transactions of the American Fisheries Society, 72(1): 97–102, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1943:PSB

[SS43b] E. V. Smith and H. S. Swingle. Percentages of survival of bluegills (Lepomis macrochirus) and largemouth black bass (Huro salmoides) when planted in new ponds. Transactions of the American Fisheries Society, 72(1):63–67, January 1943. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1945:SST

[SS45] Lloyd L. Smith, Jr. and Beatrice S. Smith. Survival of seven- to ten-inch planted trout in two Minnesota streams. Transactions of the American Fisheries Society, 73(1):108–116, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1947:SDW

[Sto47] Udell Bennett Stone. A study of the deep-water cisco fishery of Lake Ontario with particular reference to the bloater Leucichthys hoyi (Gill). Transactions of the American Fisheries Society, 74(1): 230–249, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sumner:1940:DPL

[Sum40] Francis H. Sumner. The decline of the Pyramid Lake fishery. Transactions of the American Fisheries Society, 69(1):216–224, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sumner:1948:AGS

[Sum48] F. H. Sumner. Age and growth of steelhead trout, Salmo gairdnerii Richardson, caught by sport and commercial fishermen in Tillamook County, Oregon. Transactions of the American Fisheries Society, 75(1):77–83, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1940:SMN

[Sur40] Eugene W. Surber. Scyphidia micropteri, a new protozoan parasite of largemouth and smallmouth black bass. Transactions of the American Fisheries Society, 69(1):169–175, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1941:QSF

[Sur41] Eugene W. Surber. A quantitative study of the food of the smallmouth black bass, Micropterus dolomieu, in three Eastern streams. Transactions of the American Fisheries Society, 70(1):311–334, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1943:STP

[Sur43a] Eugene W. Surber. Scyphidia tholiformis, a peritrichous protozoan found on the gills and external surfaces of Micropterus dolomieu and Micropterus salmoides. Transactions of the American Fisheries Society, 72(1):197–203, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1943:ONA

[Sur43b] Eugene W. Surber. Observations on the natural and artificial propagation of the smallmouth black bass, Micropterus dolomieu. Transactions of the American Fisheries Society, 72(1):233–245, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1945:EVF

[Sur45] Eugene W. Surber. The effects of various fertilizers on plant growths and their probable influence on the production of smallmouth black bass in hard-water ponds. Transactions of the American Fisheries Society, 73(1):377–393, January 1945. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1947:VNC

[Sur47] Eugene W. Surber. Variations in nitrogen content and fish production in smallmouth black bass ponds. Transactions of the American Fisheries Society, 74(1):338–349, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1949:RVR

[Sur49] Eugene W. Surber. Results of varying the ratio of largemouth black bass and bluegills in the stocking of experimental farm ponds. Transactions of the American Fisheries Society, 77(1):141–151, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1940:TEL

[SV40] Oliver H. Smith and John Van Oosten. Tagging experiments with lake trout, whitefish, and other species of fish from Lake Michigan. Transactions of the American Fisheries Society, 69(1):63–84, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1949:ECL

[Swi49] H. S. Swingle. Experiments with combinations of largemouth black bass, bluegills, and minnows in ponds. Transactions of the American Fisheries Society, 76(1):46–62, January 1949. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarzwell:1942:FPB

[Tar42] Clarence M. Tarzwell. Fish populations in the backwaters of Wheeler Reservoir and suggestions for their management. Transactions of the American Fisheries Society, 71(1):201–214, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarzwell:1945:PCF

[Tar45] Clarence M. Tarzwell. The possibilities of a commercial fishery in the TVA impoundments and its value in solving the sport and rough fish problems. Transactions of the American Fisheries Society, 73(1):137–157, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tester:1941:MSP

[Tes41] Albert L. Tester. A modified scale projector. Transactions of the American Fisheries Society, 70(1):98–101, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1940:SFA

[Tho40] Fred A. Thompson. Salmonoid fishes in the Argentine Andes. Transactions of the American Fisheries Society, 69(1):279–284, January 1940. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tingley:1942:CVM

[Tin42] Frank A. Tingley. Changes in volumetric measurement of salmon eggs. Transactions of the American Fisheries Society, 71(1):284– 285, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarzwell:1943:MFI

[TM43] Clarence M. Tarzwell and Lawrence F. Miller. The measurement of fishing intensity on the lower T.V.A. reservoirs. Transactions of the American Fisheries Society, 72(1):246–256, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trautman:1941:FLN

[Tra41] Milton B. Trautman. Fluctuations in lengths and numbers of certain species of fishes over a five-year period in Whitmore Lake, Michigan. Transactions of the American Fisheries Society, 70(1): 193–208, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trembley:1945:RPT

[Tre45] Gordon L. Trembley. Results from plantings of tagged trout in Spring Creek, Pennsylvania. Transactions of the American Fisheries Society, 73(1):158–172, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Triggs:1941:RFA

[Tri41] Charles W. Triggs. Relation of the Fishery Advisory Committee to the commercial and sport fisheries. Transactions of the American Fisheries Society, 70(1):70–71, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thorpe:1947:PDB

[TRW47] Lyle M. Thorpe, H. John Rayner, and Dwight A. Webster. Population depletion in brook, brown, and rainbow trout stocked in the Blackledge River, Connecticut in 1942. Transactions of the American Fisheries Society, 74(1):166–187, January 1947. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tryon:1943:ECH

[Try43] C. A. Tryon, Jr. The effect of covering hatchery troughs on the growth of cutthroat trout (Salmo clarkii). Transactions of the American Fisheries Society, 72(1):145–149, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1942:RBP

[Van42] John Van Oosten. Relationship between the plantings of fry and production of whitefish in Lake Erie. Transactions of the American Fisheries Society, 71(1):118–121, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1947:MSO

[Van47] John Van Oosten. Mortality of smelt, Osmerus mordax (Mitchill), in Lakes Huron and Michigan during the fall and winter of 1942– 1943. Transactions of the American Fisheries Society, 74(1):310– 337, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1948:TFD

[Van48] John Van Oosten. Turbidity as a factor in the decline of Great Lakes fishes with special reference to Lake Erie. Transactions of the American Fisheries Society, 75(1):281–322, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1949:DDF

[Van49] John Van Oosten. A definition of depletion of fish stocks. Transactions of the American Fisheries Society, 76(1):283–289, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1949:AGL

[VH49] John Van Oosten and Ralph Hile. Age and growth of the lake whitefish, Coregonus clupeaformis (Mitchill), in Lake Erie. Transactions of the American Fisheries Society, 77(1):178–249, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1943:PGR

[Vio43] Percy Viosca, Jr. Phenomenal growth rates of largemouth black bass in Louisiana waters. Transactions of the American Fisheries Society, 72(1):68–71, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1945:CAP

[Vio45] Percy Viosca, Jr. A critical analysis of practices in the management of warm-water fish with a view to greater food production. Transactions of the American Fisheries Society, 73(1):274–283, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanOosten:1947:ROC

[vO47] John van Oosten. Reports of officers: The cancellation of the Seventy-Fourth Annual Meeting of the American Fisheries Society. Transactions of the American Fisheries Society, 74(1):11–15, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanOosten:1948:AFS

[vO48] John van Oosten. The American Fisheries Society — its 75th anniversary. Transactions of the American Fisheries Society, 75(1): 11–15, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanOosten:1949:ORP

[vO49] John van Oosten. Opening remarks of the President. Transactions of the American Fisheries Society, 76(1):392–393, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanOosten:1942:AFP

[vOSW^{+}42] Chairman John van Oosten, Edward Schneberger, Herbert E. Warfel, E. B. Speaker, T. H. Langlois, E. L. LeCompte, George W. Bennett, James R. Simon, and H. J. Deason. The administration of fishery programs. Transactions of the American Fisheries Society, 71(1):315–331, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallace:1941:SDC

[Wal41] David H. Wallace. Sexual development of the croaker, Micropogon undulatus, and distribution of the early stages in Chesapeake Bay. Transactions of the American Fisheries Society, 70(1):475– 482, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Washburn:1948:PCC

[Was48] George N. Washburn. Propagation of the creek chub in ponds with artificial raceways. Transactions of the American Fisheries Society, 75(1):336–350, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1943:FPY

[Web43] Dwight A. Webster. Food progression in young white perch Morone americana (Gmelin) from Bantam Lake, Connecticut. Transactions of the American Fisheries Society, 72(1):136–144, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1948:RTS

[Web48] Dwight A. Webster. Relation of temperature to survival and incubation of the eggs of smallmouth bass (Micropterus dolomieu). Transactions of the American Fisheries Society, 75(1):43–47, January 1948. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warfel:1943:SOM

[WFJ43] Herbert E. Warfel, Terrence P. Frost, and Warren H. Jones. The smelt, Osmerus mordax, in Great Bay, New Hampshire. Transactions of the American Fisheries Society, 72(1):257–262, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whiteleather:1949:FCT

[Whi49] R. T. Whiteleather. Fishery conservation through education. Transactions of the American Fisheries Society, 76(1):118–120, January 1949. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1947:AEF

[WL47] Dwight A. Webster and G. Scott Little. Angling experiments by the Fly Tyers Club at Cornell University. Transactions of the American Fisheries Society, 74(1):63–70, January 1947. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1941:FOU

[Wol41] Louis E. Wolf. Further observations on ulcer disease of trout. Transactions of the American Fisheries Society, 70(1):369–381, January 1941. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodbury:1942:SMF

[Woo42] Lowell A. Woodbury. A sudden mortality of fishes accompanying a supersaturation of oxygen in Lake Waubesa, Wisconsin. Transactions of the American Fisheries Society, 71(1):112–117, January 1942. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodbury:1943:VBD

[Woo43] Lowell A. Woodbury. Vitamin B_{1} deficiency in hatchery reared rainbow trout. Transactions of the American Fisheries Society, 72(1):30–34, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wright:1945:EMF

[Wri45] Stillman Wright. The effect of moonlight on fishing success in Fish Lake, Utah. Transactions of the American Fisheries Society, 73(1): 52–58, January 1945. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williamson:1943:RPL

[WS43] Lyman O. Williamson and Edward Schneberger. The results of planting legal-sized trout in the Deerskin River, Vilas County, Wisconsin. Transactions of the American Fisheries Society, 72(1):92– 96, January 1943. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).